### v1

# TD-Orch: Scalable Load-Balancing for Distributed Systems with Applications to Graph Processing

Abstract.
In this paper, we highlight a task-data orchestration abstraction that supports a range of distributed applications, including graph processing and key-value stores. Given a batch of tasks each requesting one or more data items, where both tasks and data are distributed across multiple machines, each task must get co-located with its target data (by moving tasks and/or data) and executed. We present TD-Orch, an efficient and scalable orchestration framework featuring a simple application developer interface. TD-Orch employs a distributed push-pull technique, leveraging the bidirectional flow of both tasks and data to achieve scalable load balance across machines even under highly skewed data requests (data hot spots), with minimal communication overhead. Experimental results show that TD-Orch achieves up to $2.7\times$ speedup over existing distributed scheduling baselines.
Building on TD-Orch, we present TDO-GP, a distributed graph processing system for general graph problems, demonstrating the effectiveness of the underlying framework. We design three families of implementation techniques to fully leverage the execution flow provided by TD-Orch. Experimental results show that TDO-GP achieves an average speedup of $4.1\times$ over the best prior open-source distributed graph systems for general graph processing.
graph processing; system orchestration; load balance; distributed-memory algorithms
conference: ACM SIGPLAN Annual Symposium on Principles and Practice of Parallel Programming; January 31–February 04, 2026; Sydney, Australiajournalyear: 2026copyright: none
1. Introduction
Distributed-memory systems, including multi-socket NUMAs (lameter2013numa), supercomputers (schneider2022exascale), compute-equipped disaggregated memory (intelagilex; calciu2021rethinking), processing-in-memory (upmem; mutlu2023primer), computational storage (samsungsmartssd; gu2016biscuit) and datacenters (soni2014novel; wu2015cloud; kumar2021resource) provide tremendous aggregated computation power and memory capacity. A broad class of distributed-memory applications, including key-value stores and graph processing systems, can be abstracted into repeated task–data orchestration problems. Here, a computational task maintains some context and attempts to access one or multiple data items located somewhere in the system. The goal is to co-locate the task context with (copies of) all target data on the same machine, execute a user-defined function $f$, and optionally write the result(s) back to the original data location(s).
Ideally, application developers only need to provide the high-level functions for $f$, getget, and $wb$ in Fig. 1 for their application, and the distributed system runtime ensures that tasks are executed correctly and efficiently on their target data. In this paper, we consider batch parallel task execution, where concurrent tasks are collected into batches that are executed in parallel, targeting high throughput. Thus, as shown in Fig. 1, the input to the orchestration interface (API) is an array of task contexts in a batch, where each context includes the address(es) of the task’s target data.
A simple example is provided in Algorithm 1, where conference attendance fees are charged from attendee’s deposit accounts residing in a key-value store. In this scenario, the application developer uses the API to specify the end-to-end outcome (i.e., the fees are deducted from the proper accounts), while relying on the task-data orchestration provided by the underlying system to achieve high throughput.
function Orchestration(
    n: UINT, // Number of tasks in the batch
    tasks: array, // The array of task contexts
    f: function, // Execute using f(task, data)
    get: function, // get(pointer) -> data
    wb: function, // Writeback wb(pointer, data)
    merge_value: Merge-able operation in Sec 2.3
Figure 1. Orchestration Function API. Task and available data start (roughly) evenly divided among the machines.
Input:
accounts: Distrib. data array of (conf_name, deposit) pairs
fees[$1:n$]: Distrib. task array of (data_addr, fee) pairs
1 Function f(task, data) :
2 return task.fee IF (data.conf_name == ”PPoPP”) ELSE 0
3Function get(data_addr) :
4 return accounts[data_addr]
5Function wb(data_addr, aggr_value) :
6 accounts[data_addr].deposit -= aggr_value
7Function ChargeConferenceFee(fees, accounts) :
8   Orchestration($n$, fees, $f$, getget, wbwb, ADD)
Algorithm 1 A Charge Conference Fee example
Providing such a task-data orchestration framework is challenging.
A key difficulty lies in achieving load balance in both communication and computation while keeping overhead low, thereby maximizing resource utilization and throughput. Each task requires all its data to be co-located on the same machine, by migrating the task and/or the data. Data targeted by many machines in a batch can become contention hot spots that violate load balance.
In this paper, we propose TD-Orch, an orchestration framework that supports Fig. 1’s API, while being both theoretically grounded in scalable load balancing and practically efficient. Our key technique is distributed push-pull, which exploits the bidirectional movement of pushing tasks to data and pulling data to tasks—unlike most prior approaches that only push or pull. But judicious use of push and pull alone does not solve the hot spot problem, since neither pushing tasks to a machine holding hot data nor pulling the data to the tasks can avoid the high communication to/from that machine. Instead, TD-Orch detects hot data in each batch and employs novel (i) randomly-mapped reduction/broadcast communication trees and (ii) merge-able spillover meta-task structures. Together, these ensure (a) for low contention data items, requesting tasks’ contexts can be pushed directly to where the item resides (no tree hops), (b) for high contention data items, the set of machines needing the item are stored in a communication-bounded way, and (c) load balance and minimal communication overhead are achieved. Our evaluation of hash table workloads on 2–16 machines demonstrates a geomean speedup of $2.68\times$, $1.48\times$ and $2.70\times$ over the three baselines we study (§2.5), respectively.
To further demonstrate the power of TD-Orch, we apply TD-Orch to build a distributed graph processing system. Early distributed frameworks such as MapReduce (dean2008mapreduce) and the Pregel series (malewicz2010pregel; bu2014pregelix; han2015giraph) inspired many successors, including BLAS (bulucc2011combinatorial), Distributed GraphLab (low2012distributed), GraphMat (sundaram2015graphmat), Graphulo (gadepally2015graphulo), GraphX (gonzalez2014graphx), Mizan (khayyat2013mizan), Pegasus (kang2009pegasus), PowerGraph (gonzalez2012powergraph), PowerSwitch (xie2015sync), and Presto (venkataraman2013presto). We classify more recent frameworks developed in the past decade into two major families: (i) Graph-theoretic approaches: GeminiGraph (zhu2016gemini), Gluon (dathathri2018gluon), PowerLyra (chen2019powerlyra), FaasGraph (liu2024faasgraph), Phoenix (dathathri2019phoenix), ShenTu (lin2018shentu), and TopoX (li2019topox); and (ii) Linear-algebra-based designs: Graphite (mofrad2020graphite), GraphPad (anderson2016graphpad), GraphScope (fan2021graphscope), GraphTap (mofrad2019efficient), and LA3 (ahmad2018la3).
We introduce TDO-GP, a distributed graph processing system atop TD-Orch. TDO-GP leverages TD-Orch as the primary mechanism for orchestrating task–data flow and employs implementation techniques to fully exploit the task–data layouts provided by TD-Orch. Like TD-Orch, TDO-GP is theoretically grounded, offering work-efficient bounds for the five graph algorithms implemented in this study. In addition to directly leveraging TD-Orch, we further optimize global communication, local computation, and the coordination between system components, tailoring these improvements for graph workloads to fully exploit the execution layout provided by TD-Orch. As a result, TDO-GP achieves geomean speedups of $4.1\times$ over the best prior open-source distributed graph processing systems designed for general graph problems (zhu2016gemini; mofrad2020graphite; ahmad2018la3) and up to $15\times$ in specific datasets.
In summary, the main contributions of this paper are:
  • We highlight task–data orchestration as an important practical problem in distributed applications, advocating a particular interface (Fig. 1).
  • We design TD-Orch, a theoretically grounded and practically fast orchestration framework that achieves scalable load balance with low communication overhead.
  • We develop TDO-GP, a distributed graph processing system built on TD-Orch and enhanced with tailored implementation techniques, which significantly outperforms prior systems (GeminiGraph, Graphite, LA3).
2. The Task-Data Orchestration Problem
In this section, we introduce an orchestration framework, TD-Orch, which aligns with the API in Figure 1. It yields optimal computation and communication costs while achieving scalable load balance. We will show that our design outperforms three prior methods for distributed load-balancing both in theory and in practice.
2.1. Problem Definition
In our abstract task-data orchestration problem for distributed-memory systems, there are $P$ machines, a batch of $n$ tasks to be executed on these machines, and data to be accessed by the tasks. The tasks are executed in stages and each task has a local context holder to store intermediate results. Data are partitioned into constant-sized data chunks. At the beginning of each stage, the tasks and the data chunks are evenly divided among the machines.
Each task requires accessing $O(1)$ data chunks (could be zero) that are stored outside its local context, and the task can be completed only after all the required data chunks are accessed. Each data chunk may be required by 0 to $n$ tasks.
We allow for both tasks and data chunks to be moved between machines, and even routed through intermediate machines. Each movement of a task or a data chunk between two machines incurs a communication cost charged to both the sending and receiving machines. Executing a task incurs a computation work cost charged to the machine it executes on. Our goal is to orchestrate the tasks and data, i.e., to find an online computation scheduling and communication scheme that enables the stage to complete, such that the communication and computation work costs are (i) low and (ii) load-balanced across the machines.
Load Balance. A key challenge in achieving load balance across the machines is the contention for the data chunks, since up to $n$ tasks may require the same data chunk. We seek task orchestration that is highly effective in practice and also with formal guarantees on its effectiveness regardless of the degree of data chunk contention. For the latter, we define the following goal:
Definition 1 (Load-Balanced Stage).
A stage of tasks with total computation work $W$ and total communication $I$ is (asymptotically) load-balanced iff each of the $P$ machines incurs $O(W/P)$ computation work and $O(I/P)$ communication cost.
2.2. Prior Scheduling Strategies
Direct Pull. A naïve scheduling strategy would first eliminate duplicate requests for data chunks within each machine and then fetch all required data chunks to the corresponding tasks. This approach, commonly used in single-machine von Neumann architectures, is effective when the contention on each data chunk is low. However, data chunk contention can be high (e.g., in graph processing, where the contention may correspond to the degree of a vertex), resulting in high communication cost for machines holding “hot” data chunks and hence severe communication load imbalance.
Direct Push. Another common strategy in distributed settings is to offload each task to the machine that stores its required data chunks. However, machines holding “hot” data chunks incur high communication cost (for receiving the tasks) and high computation work (for executing the tasks it receives), resulting in severe load imbalance. Moreover, the $O(1)$ data chunks required by a task may be spread across multiple machines, in which case there is no one machine holding all its required data chunks.
Sorting-based. Prior work on the Massively Parallel Computing (MPC) model (im2023massively) has proposed theoretically-efficient solutions to the orchestration problem (e.g., (im2023massively; goodrich2011sorting)). The procedure begins by sorting tasks according to the addresses of their required data chunks. Then, MPC broadcasting (ghaffari2019massively) is used to distribute each data chunk to the corresponding tasks. After task execution, a reverse broadcasting step writes the results back to the data chunks, followed by a reverse sorting step that restores tasks to their original order. It has been shown that the computation work is load balanced, and the communication cost per machine is $\Theta(\frac{n}{p}\log_{\frac{n}{P}}{P})$ (fish2015computational), which is optimal (roughgarden2018suffles). (Note that sorting is used here solely for grouping tasks by their requested data, and hence semi-sorting would suffice.)
Although these MPC schemes are asymptotically optimal, we will show below (Section 2.5) that the constant factors are non-negligible in practice.
2.3. TD-Orch
To address the challenges of direct push or pull, we must handle potential contention. Our method, TD-Orch, exploits the two directions of data flow: push, where (the context of) a task is sent to the machine(s) storing its required data chunks, and pull, where data chunks are fetched to the machine where the task resides. Such bidirectional strategies (referred to as distributed push-pull) are rarely emphasized in prior work. Only a few prior works demonstrate that this strategy can significantly improve load balance, but they are limited to PIM settings (kang2022pimtreevldb; kang2023pimtrie; zhao2025optimal) and rely on an extremely powerful centralized coordinator (the host CPU). Our goal is to develop a more widely applicable push-pull technique that does not require a centralized coordinator.
Ideally, when contention information is available on any machine on the fly, distributed push-pull merges redundant information on one side (task/data) and dynamically chooses the flow direction toward the contended end. This dynamic decision should balance load with low communication.
Orchestration Algorithm. TD-Orch unifies information collection and push within a single substage, where the task locations for pull are also decided. Bounded-size messages are enabled for dynamic decisions, even when machines lack direct visibility into remote states. Its main algorithm to enable load-balanced task-data orchestration consists of the following four substages in each stage:
  1. (1)
    Contention Detection: The information of each task is sent to the machine storing the target data, along a communication forest (§2.3.1) from leaves to root. Contention information on each data chunk is collected during this substage, in the form of a meta-task (§2.3.2).
  2. (2)
    Task-Data Co-location: A meta-task inherently provides a highly-compressed tree for communication. Data copies are broadcast along this tree from root to leaves. (See §2.3.3.)
  3. (3)
    Task Execution: Now that all tasks and the copies of their target data are co-located on the same machine, the tasks are executed locally.
  4. (4)
    Data Write-backs: The updated value for any data modified by an executing task is written back to the original data location. (See §2.3.4.)
2.3.1. Communication Forest
During contention detection, load balance must be preserved even before the system identifies contention. This requires that each machine only sends or receives bounded-size information, while avoiding excessive communication or round overhead caused by these bounded sizes. To meet these requirements, TD-Orch adopts a tree/forest organization of machines for load-balanced communication, inspired by MPC broadcast trees (ghaffari2019massively).
A communication tree rooted at machine $i$ is a balanced $F$-ary tree with $P$ leaves, as illustrated in Figure 2. Each of the $P$ leaf nodes represents a different physical machine, while the internal nodes correspond to virtual transit nodes, each of which is mapped to one of the $P$ machines via a hash function. Notably, a physical machine can simultaneously serve as both a leaf and an internal node. The hash function is globally known to all machines.
Figure 2. An example communication tree with fanout $F=2$ in $P=8$ machines rooted at machine 11. PM stands for physical machines. VM stands for virtual transit machines, whose ID is a tuple of (root ID, BFS ordered ID). Dash lines mean that PMs do not need to send messages to themselves up the trees. In practice, $F\gg 2$.
A communication forest is defined as a set of $P$ communication trees, each rooted at a distinct machine. This structure enables efficient collection of task-related information (such as counting task occurrences) across a batch with near-optimal performance. For each task whose data resides on physical machine $i$, the current machine sends a packed message–containing the data chunk address and the task information–to its parent in the communication tree rooted at $i$. In successive communication rounds, each parent merges the information received from its children and forwards the aggregated result to its own parent in the forest. The stage completes when all information reach their respective roots, requiring $O(\log_{F}{P})$ rounds in total.
2.3.2. Meta-Task
Two additional goals are desirable within the same contention-detection substage. (i) Task contexts without contention should be directly pushed to their data locations, avoiding the need for an extra round. (ii) For data items with contended accesses, the system should also identify where the data must be pulled in the next substage (i.e., the locations of the contending tasks).
To meet these goals, we introduce an auxiliary structure called a meta-task. Meta-tasks are bounded in size to satisfy load-balance requirements. They carry uncontended tasks as they move between machines but exclude contended ones. In addition, their inherent tree structure naturally guides the broadcast routes for data in the subsequent pull substage.
As illustrated in Figure 3, a meta-task is composed of an integer level $l$ (denoted by $L_{l}$) and data contents with constant word-sizes. Each $L_{0}$ meta-task corresponds to a task in the batch and its data content holds all the local context of the task. For $L_{l}$ meta-tasks with $l\geq 1$, the data content is a physical address pointing to the memory address of some machine where a set of $L_{l-1}$ meta-tasks are stored.
A meta-task set with chunk size $C$, as shown in Figure 3, is a set structure containing meta-tasks. The number of meta-tasks inside with the same level should not exceed $C$.
A merge operation of two meta-task sets is supported. A new meta-task set with chunk size $C$ is generated, containing all meta-tasks in both of the original sets. If the number of meta-tasks with certain level $l$ in the new set exceeds $C$, all $L_{l}$ meta-tasks will be popped out of the new set and stored locally on the current machine, and then a new $L_{l+1}$ meta-task will be pushed into the set, whose data content contains a physical pointer to where these $L_{l}$ meta-tasks are stored. The merge will execute this process from the lowest level up to highest level, until the number of meta-tasks at each level is at most $C$. An example of merging two sets with chunk size $C=3$ is illustrated in Figure 4.
This merging guarantees that meta-task sets remain bounded in size. Moreover, employing multiple layers of merging ensures that broadcasting in the subsequent pull substage incurs bounded communication per machine, preventing single data from needing to broadcast to numerous children In short, the meta-task hierarchy creates a compressed, communication-efficient broadcast/reduction tree such that (unlike Figure 2) the height and size of the tree adapts to the precise amount of data contention.
Figure 3. A meta-task set with chunk size $C$, with a zoom-in into $L_{0}$ and $L_{l}(l\geq 1)$ meta-tasks contained, and the pointers.
Figure 4. Merging two meta-task sets with chunk size $C=3$. Meta-tasks on lowest level $L_{0}$ are merged first, and then from low-level $L_{1}$ throughout to the highest-level $L_{3}$.
2.3.3. Distributed Push-Pull
Using meta-tasks allow us to send both the counts of tasks that need particular data and implicit information about which set of machines possess such tasks over the communication forest. During the preprocessing, each task is wrapped in an $L_{0}$ meta-task, and meta-task sets are constructed for those tasks accessing the same data chunk. In each communication round the constructed meta-task set is sent in the message in addition to the counts of the tasks. After receiving messages in each round, each parent node merges both the task counts and the meta-task sets. Each machine that stores a data chunk will receive the final meta-task sets and implicitly knows where these tasks are stored.
Since the size of each meta-task set is bounded (though not by constant words), the meta-task collection is still nearly perfectly load balanced. Meanwhile, the hierarchical meta-task sets also provide an implicit tree spanning the subset of machines that possess related tasks, which we denote as the subset tree. When fewer than $C$ tasks target the same data chunk, these tasks are directly pushed to the corresponding machines. If more than $C$ tasks target the chunk, they must access the data copies via a broadcast over the subset tree.
2.3.4. Write Backs
If tasks require write-backs that meet a special merge-able characteristic (defined in 2), then write-backs can be performed efficiently in a bottom-up fashion using the subset trees.
Definition 2 (Merge-able Operations).
An operation $\oplus$ is merge-able if there exist operations $\odot$ and $\otimes$ such that for any data $x$ and $y_{i}(1\leq i\leq n)$, $x\oplus y_{1}\oplus...\oplus y_{n}=x\odot(y_{1}\otimes...\otimes y_{n})$.
Merge-able operations $\oplus$ are commonly seen, including and not limited to (i) idempotent operations ($\odot$ is $\oplus$ and $\otimes$ is keeping one left), (ii) set-associative operations ($\odot$ and $\otimes$ are both $\oplus$), (iii) concurrent writes where only one random write operation succeeds in a concurrent batch ($\odot$ is write and $\otimes$ is randomly select), and (iv) write operations where concurrent writes are resolved using a deterministic decision process, e.g., in a concurrent batch only the one with the smallest timestamp / transaction ID succeeds ($\odot$ is write and $\otimes$ is finding the value to write using the decision process).
Discussion. Note that TD-Orch requires only two sweeps over the communication forest, whereas MPC sorting (im2023massively; goodrich2011sorting) incurs more than three passes. This reduction highlights the potential of TD-Orch for strong practical efficiency.
2.4. Formal Guarantees
We use Bulk-Synchronous Parallel (BSP) model (valiant1990bridging; valiant1990general; valiant1988optimally) here to provide theoretical guarantees for our design. A BSP system consists of $P$ machines, each with a sequential processing element (PE) and a local memory without shared global memory. All machines are connected by a point-to-point communication network. The system operates in BSP rounds (supersteps), where periodic barrier synchronizations arer required. The BSP model accounts for the computation time (maximum computation work) and communication time (maximum communication amount) over all machines. Due to using maximum when defining, load balance is critical for achieving good performance. Readers with interests in theory can refer to Appendix A for more details of the model.
Our main Theorem 2.1 proves TD-Orch to be asymptotically optimal compared with the lower bounds in (roughgarden2018suffles). Please refer to Appendix B for its proof.
Theorem 2.1 (Main Theorem).
Assume $n=\Omega(\frac{P\log^{3}{P}}{\log\log{P}})$ and $n=O(P^{\epsilon})$ for some $\epsilon>1$. If each of the $P$ machines holds $\Theta(n/P)$ tasks at the beginning of a stage, and write backs are merge-able, then (i) TD-Orch takes $O(\frac{n}{P}\log_{\frac{n}{P}}{P})$ BSP communication time whp  [^1] We use $O(f(n))$ with high probability (whp) (in $n$) to mean $O(cf(n))$ with probability at least $1-n^{-c}$ for $c\geq{1}$. in this stage; and (ii) each machine will hold $\Theta(\frac{n}{P})$ tasks whp at the end of this stage.
2.5. TD-Orch Evaluation
Implementation. We compare four orchestration methods: direct-push, direct-pull, sorting-based, and TD-Orch. All of them use MPI for inter-machine communication. Direct-push, direct-pull, and TD-Orch are implemented directly, while the sorting-based method is built on KaDiS (axtmann2015practical; axtmann2018lightweight; axtmann2017place; harsh2019histogram), a state-of-the-art massively parallel sorting library. All the four methods follows the orchestration interface in Figure 1.
Microbenchmark. We evaluate the orchestration frameworks on a database application that uses a distributed hash table as its storage layer. Each task fetches a data item from the hash table, performs a multiply-and-add operation, and then writes the updated value back to the hash table. Contentions in the workload are handled respectively through the four distinct orchestration strategies.
We deploy the orchestration frameworks on a cluster of 16 machines, each equipped with AMD Opteron 6272 processors (64 cores, 2.1 GHz), 64 MB L3 cache, and 128 GB of memory. The machines are interconnected via a 1 Gbps network.
Figure 5 shows the results of a weak-scaling experiment with varying numbers of machines. In this setup, each machine executes 100k tasks, with target data addresses drawn from a Zipf distribution (zipf2016human) under different exponents $\gamma$. More detailed results for varying $\gamma$ are provided in Table 7.
Both TD-Orch and the sorting-based method are resilient to adversarial skew and exhibit near-linear weak scalability, but TD-Orch is faster. In multi-machine scenarios, TD-Orch achieves geometrically-averaged speedups of $2.68\times$, $1.48\times$ and $2.70\times$ over the other three baselines respectively.
(a) $\gamma=1.5$
(b) $\gamma=2.0$
Figure 5. Weak-scaling runtime (seconds) of hash-table microbenchmarks under Zipf with $\gamma=1.5$ and $\gamma=2.0$.
3. The TDO-GP Graph Processing System
We implement a distributed graph processing system called TDO-GP on top of our TD-Orch framework to demonstrate the effectiveness of TD-Orch. Following the common practice in prior vertex-centric works (zhu2016gemini; wu2021seastar; wang2023scaleg; zhu2022vctune), TDO-GP maintains two primary storage structures: one for vertices, which hold their intermediate values and are pinned to specific machines based on a predefined schema during graph ingestion; and another for edges and their weights, which can be dynamically orchestrated by the system.
Graph Preliminaries. We denote a graph by $G=(V,E)$ where $V$ is the vertex set and $E$ is the edge set. $n=|V|$ and $m=|E|$ are respectively the number of vertices and edges.
Interface. Our graph processing engine is primarily executed through a combined interface of TD-Orch (Section 2) and the EdgeMap abstraction (ShunB2013; shun2015ligraplus; dhulipala2017julienne; dhulipala18scalable).
Ligra (ShunB2013; shun2015ligraplus; dhulipala2017julienne) and GBBS (dhulipala18scalable) introduce EdgeMap, which works on a single machine and takes as input a graph $G=(V,E)$, a VertexSubset $U$ (a subset of vertices), and two boolean functions, $F$ and $C$. It applies $F$ to each edge $(u,v)\in E$ iff $u\in U$ and $C(v)=true$, returning a new VertexSubset $U^{\prime}$ consisting of all vertices $v$ for which $F(u,v)$ is applied and returns true. EdgeMap supports two execution modes—sparse and dense—ensuring that it runs in $O(\sum_{u\in U}\text{deg}(u))$ work (total computation) and $O(\log n)$ span (critical path length) on arbitrary inputs.
We refer to our new interface as DistEdgeMap, which is defined in Figure 6. We treat $G$ as a directed graph. For undirected inputs, we represent each undirected edge $\{u,v\}$ as two directed edges $(u,v)$ and $(v,u)$. $U$ is a DistVertexSubset, a distributed version of VertexSubset which represents a subset of vertices across distributed machines. An execution function $f$ is applied to all edges $(u,v)\in E$ where $u\in U$, and returns a value. All returned values are aggregated by DistEdgeMap using a user-defined merge-able operator merge_value, and the result is write_back to $v$. DistEdgeMap outputs a DistVertexSubset formed up by the vertices whose write_back returns true.
function Dist_Edge_Map(
    G=(V,E): Directed Graph,
    U: DistVertexSubset,  // Current iteration
    f: function,  // Execution
    write_back: function,
    merge_value: Merge-able operation
) -> DistVertexSubset;  // For next iteration
Figure 6. DistEdgeMap function
Example Graph Algorithms. We implement five representative graph algorithms using the DistEdgeMap in our system: breadth-first search (BFS), single-source shortest path (SSSP), betweenness centrality (BC), connected components (CC), and PageRank iteration (PR).
In all these algorithms, users only need to define the function $f$ and the write_back operation using only a few lines of code; then launch DistEdgeMap without needing to know the details of the underlying execution flow. Due to space constraints, we present only the pseudocode of the BFS algorithm in Algorithm 2 as a representative example to illustrate how TDO-GP operates. Please refer to Appendix D to see how to implement much more complicated algorithms such as BC in tens of lines of codes using our interface.
Input: $G=(V,E)$, where $|V|=n$; StartVertex
Output: Dist[$1:n$]: Distrib. vertex value array
[…]
2 round=1=1
3 Function f ( (u,v): Edge ) :
4  return round IF (Dist[$u$] == round - 1) ELSE -1
5Function write_back ( v: Vertex, aggr_value ) :
6  If Dist[$v$] == -1 and aggr_value != -1 then
7     Dist[$v$] = aggr_value
8     return True
9 return False
10Function BFS ( $G$, StartVertex ) :
11  DistVertexSubset frontier = {StartVertex}
12  Dist[StartVertex] = 0
13  While frontier is not empty do
14     frontier = DistEdgeMap ($G$, frontier, f, write_back, MAX)
15     round += 11
16 return Dist[$1:n$]
17
Algorithm 2 Breadth-First Search (BFS)
3.1. Main Design
Ingestion-Time Orchestration of Task-Data Flow. A distributed graph algorithm that can be expressed using DistEdgeMap can also be viewed as a task–data orchestration, whereas edges correspond to tasks and vertex intermediate values correspond to data. A key observation is that, regardless of the starting vertex, in each DistEdgeMap, the contention for data (vertex value) access induced by tasks (edges) is proportional to the degree of each vertex. Since vertex degrees are intrinsic to the structure of the graph, such contention is inevitable in all DistEdgeMap operations, regardless of the choice of starting vertex in an algorithm.
Given this observation, TD-Orch can be applied once as a preprocessing phase when the graph is ingested, which is sufficient to resolve all future skew in the execution of arbitrary DistEdgeMap operations. This one-time load balancing is particularly beneficial for highly skewed graphs, such as commonly encountered power-law graphs (gonzalez2012powergraph).
The preprocessing process consists of two stages. In the first stage, we begin by randomly placing each edge on an arbitrary machine and then execute one round of TD-Orch, which attempts to colocate all edges with their source vertices. Edges whose sources are low-degree vertices are now placed on the same machine. For edges with high-degree sources, a tree structure of meta-tasks (denoted as source trees) is created to represent their source-side communication. We then fix the storage location of these edges and propagate future source-vertex values through the source trees. In the second stage, another round of TD-Orch is performed, where a task is assigned to each edge to reach its destination vertices.
This again results in a tree structure of meta-tasks (denoted as destination trees), this time representing the paths along which the write_backwrite\_back values to the destination vertices will be sent in a DistEdgeMap.
Adopting such a preprocessing provides two key benefits: (i) it organizes edge storage into chunks, ensuring balanced computation load; and (ii) it establishes a tree-shaped communication framework for both get_value and write_back, ensuring balanced communication with reduced overhead.
Sparse-Dense Execution. Ligra (ShunB2013) and GBBS (dhulipala18scalable) introduces single-machine sparse-dense execution. We design a new dual-mode execution flow for distributed graph processing which can switch between sparse and dense execution modes to efficiently exploit the data layouts provided by TD-Orch.
The mode is determined by the size of the current DistVertexSubset. Specifically, the sparse mode is optimized for small DistVertexSubset, while the dense mode is optimized for large ones.
In the sparse mode, a vertex-centric execution flow is used: each active vertex $u\in U$ propagates its value down the source tree, activating all corresponding edges. The result values are then sent and merged along the destination tree.
In contrast, the dense mode uses an edge-centric execution flow. All active vertices $u\in U$ broadcast their values to all machines. Each edge $(u,v)$ is then executed if $u\in U$.
The choice between sparse and dense mode is dynamically determined by the size of the active frontier in the current round. Specifically, if $\Sigma_{u\in{U}}deg(u)=o(P\cdot|U|)$ where $d(\cdot)$ denotes the out-degree of inter-machine edges, the sparse mode is used; otherwise, the dense mode is employed.
This dynamic switching ensures that the execution remains both work- and communication-efficient in theory (with work and communication of $O(\Sigma_{u\in{U}}deg(u))$ as in Ligra and load-balanced due to the use of TD-Orch), while also achieving high performance in practice.
3.2. Implementation Techniques
In TDO-GP, we employ several implementation techniques, which can be grouped into three categories. Due to space limit, we summarize only the key functionalities of each group here. Detailed descriptions are provided in Appendix E.
3.2.1. Optimized Global Communication
We adopt deduplication and destination-aware broadcasting to reduce inter-machine communication, alleviating this critical bottleneck.
3.2.2. Work-Efficient Local Computation
Our design leverages ParlayLib (blelloch2020parlaylib), a state-of-the-art parallel library, to implement work-efficient computations. We enhance VertexSubset with a concurrent hash table and a concurrent bitmap, and we further develop a lightweight implementation of local EdgeMap. Moreover, DistEdgeMap accepts an optional filter_dst function, excluding vertices that are guaranteed not to be destinations in the current round.
3.2.3. Aligned Coordination between Components
We design tighter coordination between system components—between threads and scheduler, among threads to prevent cache thrashing, and between local and global execution.
3.3. Formal Guarantees
Since Theorem 2.1 applies to DistEdgeMap, all of our algorithms are work-efficient and load-balanced across machines. The theoretical bounds summarized in Table 1 are achieved when $n=\Omega(P\log^{2}{P})$. While most prior works (zhu2016gemini; mofrad2020graphite; ahmad2018la3; yan2015effective; lu2014large; yan2014pregel) mostly focus on empirically improve their systems, we place strong emphasis on graph-theoretic guarantees as a guidance for system implementation.
As demonstrated in Section 4.2, this theoretical grounding directly translates into practical performance improvements.
Table 1. Theoretical bounds of graph algorithms.
Alg. BFS SSSP BC $C$ PR
Comp. $O\left(\frac{m}{P}\right)$ $O\left(\frac{m}{P}\right)$ $O\left(\frac{m}{P}\right)$ $O\left(\frac{m}{P}\cdot{\text{diam}}\right)$ $O\left(\frac{m}{P}\right)$
Comm. All equal to $O(\text{Computation}\cdot{\log_{\frac{n}{P}}{P}})$
4. Evaluation
4.1. Experimental Setup
Platform. We evaluate all methods on a cluster of 16 machines, each equipped with AMD Opteron 6272 processors (64 cores, 2.1 GHz), 64 MB L3 cache, and 128 GB of memory. The machines are interconnected via a 10 Gbps network.
Competitors. Due to their strong reported performance, we select GeminiGraph, Graphite, and LA3 as baselines. GeminiGraph (zhu2016gemini) is a computation-centric distributed framework that employs a hybrid model and cache-aware optimizations to reduce communication and improve scalability. Graphite (mofrad2020graphite) is a NUMA-aware graph analytics system using thread-based partitioning and diagonal scaling to minimize communication costs while supporting linear-algebra-based graph computations. LA3 (ahmad2018la3) is a hybrid system combining a vertex-centric interface with a linear algebra engine featuring optimizations in exploiting sparsity. Additionally, we include Pregel+ (yan2015effective; lu2014large; yan2014pregel) in one end-to-end experiment in §4.2 and GBBS (dhulipala18scalable) in the ablation study presented in §4.5.
Datasets. We evaluate all methods on Reddit (hamilton2017inductive), uk-2005 (BoVWFI; BRSLLP), Twitter-2010 (kwak2010twitter; boldi2011layered), friendster (yang2012defining), Hyperlink-2012 (meusel2015graph) and Road-USA (road).
See the first column of Table 2.
4.2. End-to-End Results
Table 2 presents a performance comparison between our proposed framework, TDO-GP, and three state-of-the-art systems: Gemini (zhu2016gemini), Graphite (mofrad2020graphite), and LA3 (ahmad2018la3), evaluated on a diverse set of graph datasets. We select the number of machines roughly proportional to the number of vertices in each dataset, except for particularly demanding datasets (e.g., Road-usa) that require excessive runtime. Overall, TDO-GP consistently achieves superior performance, outperforming the baselines in most scenarios often by substantial margins.
For social media graphs such as Twitter-2010, Friendster, and Reddit, TDO-GP achieves speedups more than $2\times$ in $19$ out of the $30$ tests we carry out. These gains stem from the ability of TDO-GP to more effectively exploit graph sparsity while minimizing inter-machine communication overhead.
For large-diameter graph datasets, such as Hyperlink-2012 (web graph), uk-2005 (web graph), and Road-USA (road network), the performance gains of TDO-GP are even more pronounced. The most striking example is Road-USA, where TDO-GP achieves more than $15\times$ speedup over Gemini and over $100\times$ compared to Graphite and LA3. These improvements arise from the complexity of $O(n\cdot\text{diam}+m)$ work bounds in the baseline algorithms, compared with the work bounds of $O(n+m)$ in TDO-GP. In high-diameter graphs, where few vertices are active per round but the number of rounds is large, baseline inefficiencies are amplified, whereas TDO-GP maintains stable and efficient performance.
It is worth noting that TDO-GP underperforms the baselines in only $2$ cases out of the $30$ experiments: Twitter-2010 PR and Road-USA PR. Detailed explanations for these exceptions are provided in Section 4.5.
We also compared TDO-GP against Pregel+ (yan2015effective; lu2014large; yan2014pregel). However, their Hadoop-based implementations are at least $5\times$ slower than all the other baselines considered here and are therefore omitted from subsequent analyses.
Table 2. End-to-end runtime comparison (in seconds) across various datasets and algorithms.
Dataset Alg. TDO-GP Gemini Graphite LA3
Reddit BFS 0.015 0.151 0.130 0.157
4 machines SSSP 0.071 0.172 0.141 0.505
n=2.33n=2.33M BC 0.076 0.275 0.507 0.437
m=114m=114M $C$ 0.016 0.211 0.266 0.663
diam [^2] Computing the exact diameter of large graphs is often prohibitively expensive. In our results, an asterisk (*) denotes an estimated diameter obtained using the approximation method in Ligra (ShunB2013), whereas diameters without an asterisk correspond to values reported by the dataset providers. $6^{*}$ PR 0.143 0.676 0.297 1.92
uk-2005 BFS 0.759 8.98 8.01 32.56
8 machines SSSP 0.820 10.68 13.79 45.41
n=39.5n=39.5M BC 3.00 20.23 34.79 75.90
m=482m=482M $C$ 3.30 11.22 20.85 63.11
diam $276^{*}$ PR 2.37 2.77 7.69 9.06
Twitter2010 BFS 0.938 1.35 3.13 7.67
8 machines SSSP 2.05 3.63 3.36 7.94
n=41.7n=41.7M BC 3.04 4.34 6.16 15.54
m=1.47m=1.47B $C$ 1.28 7.65 7.40 21.73
diam $23^{*}$ PR 10.46 12.72 9.85 13.67
friendster BFS 1.98 3.12 3.70 12.38
8 machines SSSP 3.09 4.42 4.25 27.23
n=65.6n=65.6M BC 5.29 8.24 7.81 69.15
m=1.80m=1.80B $C$ 2.52 17.92 9.92 56.25
diam =32=32 PR 18.32 33.07 21.71 28.03
Hyperlink12 BFS 1.60 6.22 5.59 20.50
16 machines SSSP 2.40 8.02 28.53 33.10
n=102n=102M BC 3.03 16.09 46.15 51.50
m=0.93m=0.93B $C$ 5.26 11.71 29.42 48.31
diam $95^{*}$ PR 7.12 10.10 14.43 38.78
Road-USA BFS 10.68 357.73 1224.83 1431.11
16 machines SSSP 15.76 362.11 1651.52 1788.53
n=23.9n=23.9M BC 26.24 763.09 >>35min >>35min
m=28.9m=28.9M $C$ 12.93 383.51 1814.93 >>35min
diam $6139^{*}$ PR 1.00 0.612 3.01 3.59
(a) SSSP on Twitter.
(b) BC on Twitter.
(c) PR on Twitter.
Figure 7. Strong scalability on Twitter with 1–16 machines.
4.3. Scalability
Strong Scalability. We run the SSSP, BC, and PR algorithms on the Twitter dataset using a different setup of 1, 2, 4, 8, and 16 machines. The results in Figure 7 demonstrate that TDO-GP exhibits near linear scalability across all machine configurations, while other methods either show poor scalability when increasing the number of machines, or exhibit poor performance. Similar to Table 2, the only case we are not winning is the PR against Graphite, where further details on this behavior will be discussed in Section 4.5.
Weak Scalability. We also evaluate weak scalability on graph systems by varying the number of machines while keeping the graph size per machine fixed. To construct graphs in such scale, we employ two graph generation methods: Erdős-Rényi (ER) (erd6s1960evolution), which produces unskewed graphs, and Barabási–Albert (BA) (barabasi1999emergence), which produces skewed graphs. The average number of edges per machine is fixed at 40M for a fair comparison, where in the 16-machine case the graph is in similar size as uk-2005. For the BA model, we set the power-law exponent to $\gamma=2.2$, consistent with the measured skew observed in natural graphs reported by PowerGraph (gonzalez2012powergraph). We run both PR and BC on $P=1,2,4,8,16$ machines. In Figure 9, the x-axis represents the number of machines and the y-axis represents the execution time. The results show that TDO-GP maintains nearly constant runtime across different machine counts, demonstrating excellent weak scalability, whereas prior works exhibit significant scalability degradation.
(a) ER-PR
(b) ER-BC
(c) BA-PR
(d) BA-BC
Figure 8. Weak scalability results. $y$-axis is runtime (in seconds) and $x$-axis is the number of machines used.
4.4. Ablation Study
Execution Time Breakdown. As shown in Figure 9, on 16 machines the time spent on pure local computation is effectively reduced across all algorithms. At the same time, the communication time (mainly for propagating vertex values and the cost of synchronization rounds) and the overhead time (including data preparation and data receiving for multi-machine transmission) are also non-negligible, and together contribute significantly to the overall execution time.
Figure 9. Runtime on Twitter with 16 machines. Breakdown into Communication, Computation, and Overhead.
Sensitivity to Each Optimization. We first demonstrate the effectiveness of TD-Orch within TDO-GP by comparing its final design against a baseline of Ligra with direct-pull, which performs no load balancing across machines. As shown in Table 3, our BC results on the Twitter graph achieve up to $220\times$ speedup.
Additionally, we evaluate the sensitivity of TDO-GP to the implementation techniques introduced in Section 3.2. Using a controlled-variable approach, we measure the slowdown incurred when removing each optimization family from the fully optimized TDO-GP. Table 4 reports the results on the Twitter dataset using 4, 8, and 16 machines. The results show that disabling any single optimization family causes substantial performance degradation, with slowdowns of up to $5.72\times$, underscoring the critical importance of each technique in achieving high practical efficiency.
Table 3. Runtime (in seconds) of BC on Twitter.
Machine Num. 1 4 8 16
Ligra-Dist 5.36 563.11 670.71 321.60
TDO-GP 4.54 3.70 3.05 2.35
Table 4. Slowdown when removing individual techniques (see Section 3.2) from the fully optimized TDO-GP.
Number of Machines 4 8 16
Final Optimized TDO-GP (Runtime in seconds).
BFS: TDO-GP 1.04 0.94 0.87
SSSP: TDO-GP 2.65 2.06 1.54
BC: TDO-GP 3.70 3.05 2.35
PR: TDO-GP 15.80 11.91 9.03
Removing 3.2.1: Optimized Global Communication.
[…]
Removing 3.2.2: Work-Efficient Local Computation.
[…]
Removing 3.2.3: Aligned Coordination.
[…]
4.5. Ablation Study on NUMA Effects
As discussed in Sections 4.2 and 4.3, the only case where TDO-GP does not outperform prior systems is PageRank (PR), and only against Graphite (mofrad2020graphite). TDO-GP is built upon ParlayLib (blelloch2020parlaylib), a fast parallel library that achieves strong scalability with respect to the total number of cores, but is NUMA-oblivious. While this design works well on machines with high-speed NUMA interconnects, our 16-machine cluster has budget constraints: each machine contains four NUMA nodes connected in a square topology, where certain NUMA-to-NUMA accesses are significantly slower. This architectural limitation hinders the performance of TDO-GP on PR.
We conduct two additional ablation studies to evaluate the effectiveness of TDO-GP when NUMA-related limitations are mitigated. First, we perform a multi-machine experiment on the same cluster used in Section 4.2, but restrict execution to a single NUMA node per machine. The results, shown in Table 5, demonstrate that the performance gap between TDO-GP and baselines is substantially reduced. Moreover, TDO-GP achieves a significantly smaller memory footprint compared to Graphite. The remaining minor performance difference arises because ParlayLib (blelloch2020parlaylib), the underlying parallel library of TDO-GP, is optimized for many-core settings, making the single-NUMA configuration less favorable.
We further conduct a single-machine experiment on a more powerful server with superior NUMA interconnects. This machine is equipped with four Intel(R) Xeon(R) E7-8867 v4 @2.40GHz processors, providing a total of 144 cores and 180MB of L3 cache. Unlike our 16-machine cluster, it features an all-to-all interconnection across its four NUMA nodes, eliminating the bottlenecks of square-topology interconnects. The results on the Twitter dataset are presented in Table 6, where TDO-GP consistently outperforms all other baselines by a significant margin. Notably, TDO-GP even surpasses GBBS (dhulipala18scalable), the state-of-the-art single-machine implementation, owing to our optimized lightweight local EdgeMap.
Given the results of these ablation studies, we are confident that TDO-GP can be efficiently deployed to address all the graph problems in this paper on clusters with fast NUMA interconnections, which are typical in high-performance computing environments and modern datacenters.
Table 5. Runtime (in seconds) of PR on Twitter. Each machine uses only 11 NUMA node. OOM is out-of-memory.
Machine Num. 1 4 8 16
Gemini 46.67 30.72 25.61 18.59
Graphite OOM OOM 22.18 14.69
TDO-GP 40.74 32.30 21.67 15.07
Table 6. Twitter runtime (in seconds) on Xeon E7-8867 v4.
Method BFS BC PR
Gemini 0.29 1.17 4.72
Graphite 2.17 3.91 10.05
GBBS 0.29 1.05 6.81
TDO-GP 0.12 0.95 4.43
5. Conclusion
In this paper, we introduce an abstraction of the task–data orchestration problem, unifying a range of distributed applications. We design TD-Orch to realize this orchestration with scalable load balance and low communication. Building on TD-Orch, we develop TDO-GP, a distributed graph processing system enhanced with implementation techniques for graph workloads. Both TD-Orch and TDO-GP are theoretically grounded and achieve high practical performance.
Acknowledgements.
This research was supported by the Parallel Data Lab (PDL) Consortium (Amazon, Bloomberg, Datadog, Google, Honda, Intel, Jane Street, LayerZero, Meta, Microsoft, Oracle, Pure Storage, Salesforce, Samsung, and Western Digital).
Appendix A Models for Distributed Computing
In this paper, we model and analyze algorithms using the classic Bulk-Synchronous Parallel (BSP) model (valiant1990bridging; valiant1990general; valiant1988optimally). The BSP consists of $P$ machines, each with a sequential processing element (PE) and a local memory. There is no shared global memory. All machines are connected by a point-to-point communication network. All remote accesses to data on other machines are carried out via messages sent over this network. The system operates in bulk-synchronous rounds (supersteps), where periodic barrier synchronizations separate periods of asynchronous computation and communication.
Cost Metrics. The BSP model defines the notion of an $h$-relation of a superstep, which combines the maximum computation work, maximum communication amount and synchronization time over all machines. Due to using maximum when defining $h$-relations, load balance is critical for achieving good performance. In all our analysis, for clarity, we decouple the computation time and communication time, and do not consider the BSP parameters $g$ and $L$ in (valiant1990bridging; gerbessiotis1994direct). Computation time $t$ and communication time $h$ can be combined into a formal BSP result $\Theta(gh+t)$ as long as the synchronization overhead can be hidden by system communication, i.e., $(gh+t)\in\Omega(L)$.
Other Distributed Models. The Massively Parallel Model of Computation (MPC) (im2023massively; karloff2010model; goodrich2011sorting; beame2013communication; andoni2014parallel) family of models regards synchronization as the main bottleneck in distributed systems and takes communication rounds as the only optimization target. We choose the the BSP model because it differs from MPC in two ways: (i) The the BSP model accounts for the total communication among machines rather than ignoring the amount of communication in each round. (ii) The the BSP model also considers computation costs neglected by the MPC model, which turn out to be critical for real-world system designs on emerging distributed hardware (kang2022pimtreevldb; bindschaedler2020hailstorm) other than MapReduce systems  (dean2008mapreduce).
LOCAL/CONGEST models (linial1987distributive; linial1992locality) are a popular model family for distributed systems. Like the models we consider, each machine has only a local view of the computation. These models consider network topology, which we ignore in the BSP model. However, they have similar weaknesses to the MPC model with regards to their metrics: (i) each machine can send/receive up to its maximum allowed communication at each round, whereas we seek to minimize communication, and (ii) computation costs are not considered.
The Processing-in-Memory (PIM) model (kang2021processing) is a recently proposed model for emerging compute-in-memory hardware (mutlu2023primer). The model has a powerful Host CPU and $P$ PIM nodes, each comprised of a PE and a memory. It operates in bulk-synchronous rounds, but unlike the BSP model, all communication must be sent by or to the Host CPU. Furthermore, the model decouples the computation time of the CPU from other processors, resulting in a lack of analysis of load-balance between them.
Appendix B Proof for Theorem 2.1
We will use the balls-into-bins game of Lemma B.1 to prove load balance.
Lemma B.1 (Weighted Balls into Bins  (sanders1996competitive; nievergelt1973binary)).
Placing weighted balls with total weight $W=\Sigma w_{i}$ and weight limit $W/(P\log P)$ into $P$ bins uniformly randomly yields $\Theta(W/P)$ weight in each bin whp.
Proof for Theorem 2.1.
We set $C=\Theta(1)$ and $F=\Theta(\frac{n}{P\log^{2}{P}})$. (i) Substage 1 takes at most $O(\log_{F}{P})$ rounds and in each round at most $O(n)$ words of messages are sent. Since meta-task sizes are bounded by $O(C\log_{C}{n})\leq{O}(\log{P})$, the weight limit in Lemma B.1 is $O(FC\log_{C}{n})\leq{O}(\frac{n}{P\log{P}})$ and the communication is thus load-balanced. When $n=\Omega(\frac{P\log^{3}{P}}{\log\log{P}})$, $\Theta(\log{F})=\Theta(\log{\frac{n}{P}})$, so the communication time is $O(\frac{n}{P}\log_{\frac{n}{P}}{P})$ in substage 1. Substage 2 takes less communication time than substage 1 because it only requires partial communication over substage 1’s trace, so (i) is proved. (ii) Since meta-tasks have size $O(C\log_{C}{n})$, Lemma B.1 can also be applied. ∎
Appendix C Detailed Orchestration Result
Table 7. Runtime (in seconds) of a weak-scaling microbenchmark of hash table workloads with different numbers of machines under different Zipf parameters $\gamma$.
Workload 2 4 8 16
$\gamma=1.5$
Direct Push 0.1093 0.1747 0.4696 0.5869
Direct Pull 0.0933 0.1572 0.2818 0.3542
Sorting 0.2545 0.3044 0.5397 0.5365
TD-Orch 0.0895 0.1082 0.1590 0.1863
$\gamma=2.0$
Direct Push 0.1697 0.2770 0.5072 0.9655
Direct Pull 0.0903 0.1609 0.2536 0.4579
Sorting 0.4499 0.3080 0.3717 0.4349
TD-Orch 0.0847 0.1280 0.1468 0.1392
$\gamma=2.5$
Direct Push 0.1480 0.3024 0.8190 1.5284
Direct Pull 0.0939 0.1197 0.3183 0.5991
Sorting 0.2030 0.2544 0.5029 0.5771
TD-Orch 0.0936 0.1174 0.2417 0.2937
Appendix D Algorithm for Betweenness Centrality
Algorithm 3 describes the betweenness centrality (BC) computation from a single root vertex, which is commonly used in performance tests in prior works (ShunB2013; dhulipala18scalable; zhu2016gemini).
Input: $G=(V,E)$; StartVertex
Output: BCValues: distributed vertex value array
1 Initialize distributed vertex value array: NumPaths[$1:n$], BCValues[$1:n$], rounds[$1:n$]
2 round=1=1
3 Function f_forward ( (u,v): Edge ) :
4 return NumPaths[$u$]
5Function write_back_forward ( v: Vertex, aggr_value ) :
6 If rounds[$v$] == 0 or rounds[$v$] == round then
7    NumPaths[$v$].fetch_add(aggr_value)
8    If rounds[$v$] == 0 then
9       rounds[$v$] = round
10       return True
11 return False
12Function f_backward ( (u,v): Edge ) :
13 return BCValues[$u$]
14Function write_back_backward ( v: Vertex, aggr_value ) :
15 If rounds[$v$] == round then
16    BCValues[$v$].fetch_add(aggr_value)
17 return False
18Function BC ( $G$, StartVertex ) :
19 DistVertexSubset frontier = {StartVertex}
20 frontiers = {frontier}
21 NumPaths[StartVertex] = 1
22 rounds[StartVertex] = 1
 // Forward Pass
23 While frontier is not empty do
24    round += 11
25    frontier = DistEdgeMap ($G$, frontier, f_forward, wb_forward, add)
26    frontiers.append(frontier)
 // Backward Pass
27 parallel for $i\in[1:n]$ do
28    BCValues[$i$] = NumPaths[$i$] = 1/1/NumPaths[$i$]
29 While round >0>0 do
30    frontier = frontiers[round]
31    DistEdgeMap ($G$, frontier, f_backward, wb_backward, add)
32    round -= 11
33
34 parallel for $i\in[1:n]$ do
35    BCValues[$i$] = BCValues[$i$] / NumPaths[$i$] - 1
36 return BCValues[$1:n$]
37
Algorithm 3 Betweenness Centrality (BC)
Appendix E Details for Implementation Techniques
We present here the implementation techniques used in our graph processing system, which efficiently exploit the data layout and execution flow provided by TD-Orch, leading to the strong performance demonstrated in Section 4.
E.1. Optimized Global Communication
As we will show in Section 4.4, the critical bottleneck in a distributed graph system is the inter-machine communication time. Optimizing global communication is therefore essential. First, we apply deduplication during message passing along the meta-task trees, aggressively merging all repeated items. Second, in dense mode, we reduce broadcast traffic by sending the values of each vertex only to the machines that store its corresponding edges. The mapping of vertices to machines is established during the preprocessing phase.
E.2. Work-Efficient Local Computation
For all local computations in our implementation, we employ work-efficient designs that are both theoretically optimal and practically fast. Our design leverages ParlayLib (blelloch2020parlaylib), a state-of-the-art C++ parallel library providing a wide range of work-efficient parallel algorithms as primitives.
The DistVertexSubset structure consists of a collection of VertexSubset instances, one per machine, where each VertexSubset can independently switch between two representation formats introduced in Ligra (ShunB2013). We introduce two implementation improvements. (i) In the sparse representation, we replace their array of vertices with a phase-concurrent hash table with adaptive sizing, enabling more efficient lookups within VertexSubset. (ii) In the dense representation, we replace the original parallel C++ Boolean-map with a concurrent bitmap, improving cache efficiency.
For local edges whose source and destination vertices reside on the same machine, we bypass the global orchestration framework and instead execute a local EdgeMap. We implement a lightweight EdgeMap entirely using ParlayLib primitives, which, as we show in Section 4.5, is more efficient than the ones in Ligra/GBBS for EdgeMap workloads.
Additionally, we allow users to provide an optional input function filter_dst to DistEdgeMap to accelerate execution by filtering out vertices that are guaranteed not to be destination vertices in the current round. Disabling filter_dst does not affect correctness but may reduce efficiency.
E.3. Aligned Coordination between Components
A distributed graph processing system involves multiple software components running across diverse hardware resources. Ensuring a coordinated execution schema among these components is crucial. Our implementation optimizes this coordination to achieve aligned and efficient execution.
For better alignment with the parallel thread scheduler in ParlayLib, we avoid using nested ParallelFor, which is commonly seen in graph algorithms but inefficient under such schedulers. Instead, we adopt Flattern, assisted by delayed structures (blelloch2020parlaylib), to transform two-dimensional nested parallelism into one-dimensional parallelism.
To better align local execution across threads and minimize cache thrashing, we design concurrent counters and other shared concurrent structures with sizes padded to a multiple of the cache-line size. This padding prevents false sharing, ensuring that accesses from different cores do not contend on the same cache line.
We also align local computation with the load-balanced data layout across machines. We initialize the graph with a vertex-value layout such that the total number of outgoing edges (out-degrees) assigned to each machine is approximately equal. This ensures that most local computations are naturally load-balanced across machines, since each handles a comparable number of edges.