### v2

# TD-Orch: Scalable Load-Balancing for Distributed Systems with Applications to Graph Processing

Abstract
In this paper, we introduce a task-data orchestration abstraction that supports a range of distributed applications, including graph processing and key-value stores. Given a batch of lambda tasks each requesting one or more data items, where both tasks and data are distributed across multiple machines, each task must be co-located with its target data (by moving tasks and/or data) and then executed. We present TD-Orch, an efficient and scalable orchestration framework featuring a simple application developer interface. TD-Orch employs a distributed push-pull technique, leveraging the bidirectional flow of both tasks and data to achieve scalable load balance across machines even under highly skewed data requests (data hot spots), with minimal communication overhead. Experimental results show that TD-Orch achieves up to $2.8\times$ speedup over existing distributed scheduling baselines.
Building on TD-Orch, we present TDO-GP, a distributed graph processing system for general graph problems, demonstrating the effectiveness of the underlying framework. We design three families of implementation techniques to fully leverage the execution flow provided by TD-Orch. Experimental results show that TDO-GP achieves an average speedup of $4.1\times$ over the best prior open-source distributed graph systems for general graph processing.
1 Introduction
Distributed-memory computer systems, including supercomputers [schneider2022exascale], multi-socket NUMAs [lameter2013numa], compute-equipped disaggregated memory [intelagilex, calciu2021rethinking], processing-in-memory [upmem, mutlu2023primer], computational storage [samsungsmartssd, gu2016biscuit], and datacenters [soni2014novel, wu2015cloud, kumar2021resource] provide tremendous aggregated compute power and memory capacity. Motivated by the promise of these platforms, there has been a huge body of work over the past two decades on building high-performance distributed systems for applications such as graph processing [zhu2016gemini, chen2019powerlyra, gonzalez2014graphx, gonzalez2012powergraph, mofrad2020graphite], machine learning [moritz2018ray, narayanan2019pipedream, thorpe2021dorylus, abadi2016tensorflow, chen2015mxnet], key-value store [decandia2007dynamo, escriva2012hyperdex, chang2008bigtable, cooper2008pnuts], database [corbett2013spanner, shute2013f1, dageville2016snowflake, kim2025nocap], among many other application areas.
However, writing correct and high-performance applications on distributed-memory systems is notoriously challenging. There are mainly three strategies to implement distributed applications: (i) Use high-level interfaces such as Spark [zaharia2016apache], which are easy to program but leave notable performance on the table, since they are not optimized for specific end-to-end applications. (ii) Write low-level code using interfaces like the message passing interface (MPI), which introduces significant complexity in programming and system management. (iii) Use application-specific frameworks such as PowerGraph [gonzalez2012powergraph], which are optimized for certain domains but do not support new applications outside their scope. Consequently, these approaches often lead to wasted effort and repeated reinvention of techniques already embedded in other systems. For example, many graph frameworks repeatedly reimplement the ghost/mirror node design [gonzalez2012powergraph, zhu2016gemini, chen2019powerlyra, dathathri2019phoenix].
Ideally, one could leverage a high-level programming abstraction that enables users to write correct and high-performance distributed-memory applications while avoiding the complexity typically associated with writing high-performance distributed systems by hand, or the limitations of domain specific interfaces. For example, users could specify a description of the computations performed by the application while separating data from computational logic, e.g., by specifying: (i) which data to read from the system, (ii) the computation to perform on the retrieved values, and (iii) which results to write back. Such an abstraction crucially relies on an underlying system framework that automatically and efficiently executes the described application on the distributed system, handling all the underlying orchestration of computation and communication.
In particular, a common challenge of such orchestration is how to achieve load balance in both computation and inter-machine communication while minimizing overhead. Prior work tackles this in an ad-hoc fashion, e.g., through ideas like the above-mentioned mirror/ghost nodes in graph processing [zhu2016gemini, gonzalez2012powergraph], virtual servers in key-value stores [stoica2003chord, decandia2007dynamo], and workload-aware partitioning in databases [curino2010schism, quamar2013sword]. Despite differing implementations, the high-level ideas among these techniques are similar, suggesting that a clean abstraction could unify these approaches under a shared underlying system.
In this work, we build upon prior work on *-centric (e.g., node-centric, edge-centric, data-centric) abstractions by introducing a Task-Data Orchestration abstraction based on microservices-inspired lambda-tasks. Our lambda-task-centric abstraction can efficiently support a wide range of distributed applications, including graph processing, key-value stores, databases, failure-tolerant storage and others mentioned in §2.1. In our abstraction, users or application developers only need to specify the basic operation(s) to be performed by each lambda-task. The underlying system automates efficient orchestration of computation and communication, even under workload skew.
Abstraction. Our task-data orchestration abstracts any application that fits the following pattern. A batch of lambda-tasks (referred to as simply tasks in the rest of the paper) runs in parallel, where each task resembles a C++ closure, containing local metadata (called the task context) and a local lambda function representing the required computation, along with the locations of the inputs and outputs of the closure. The application proceeds in repeated bulk-synchronous parallel (BSP) [valiant1990bridging, valiant1990general, valiant1988optimally] orchestration stages, and in each stage each task (i) reads input data from outside its context, (ii) performs a local lambda computation using its context and the read values, and (iii) writes outputs to external locations.
As a concrete example, consider a distributed hash table that serves batches of search and update operations. Here, the tasks correspond to operations and the data reside in the hash-table storage. Each batch can be executed using a single orchestration stage that: (i) reads values from the hash table using task-specified keys, (ii) processes fetched values, and (iii) writes results back to the hash table.
Another concrete example is breadth-first search (BFS) for graphs. The algorithm proceeds in synchronous stages, maintaining a frontier of vertices $i$ steps away from the source during the $i$-th stage. Here, the tasks correspond to edges, and the data correspond to the distance values of vertices. During an orchestration stage, each edge (task) reads the distance values of its endpoints (data), checks whether one endpoint lies in the current frontier, and, if so, updates the distance of the other endpoint. This also fits into the orchestration abstraction of read–execute–write. A formal example of a BFS using this interface is shown in Algorithm˜1.
function Orchestration(
tasks: An array of Task structures
wb: function, // wb(pointers[], values[])
struct Task {
pointer InputPointers[];
pointer OutputPointers[];
LocalContexts[]; // Per-Task local metadata
Lambda f; // f(InValues[]) -> OutValues[]
Figure 1: The Task-Data Orchestration interface. Task and data are distributed among the same set of machines.
More formally, a single orchestration stage follows the interface in Fig. 1. Each task resembles a C++ closure: it contains some LocalContexts, a set of InputPointers indicating which external data items it will read, a local lambda function $f$ describing the per-task computation, and a set of OutputPointers indicating where returned results will be written. The lambda $f$ takes as input the local contexts and the values read from the InputPointers, and produces return values that will be written back to OutputPointers.
System Design. In this paper, we propose TD-Orch, a new system that efficiently supports Fig. 1’s API, while being both theoretically grounded in scalable load balancing and practically efficient. Our key technique is distributed push-pull, which exploits the bidirectional movement of pushing tasks to data and pulling data to tasks—unlike most prior approaches that only push or pull. Judicious use of push and pull alone does not solve the hot spot problem, since neither pushing tasks to a machine holding hot data nor pulling the data to the tasks can avoid the high communication to/from that machine. Instead, TD-Orch detects hot data in each batch and employs novel (i) transit machines based on randomly-mapped communication trees and (ii) merge-able meta-task structures. Together, these ensure (a) for low contention data items, requesting tasks’ contexts can be pushed directly to where the item resides (no hops on a communication tree), (b) for high contention data items, the data will be sent to the tasks in a communication-efficient way, and (c) load balance and minimal communication overhead are achieved. Our evaluation of hash tables on 2–16 machines demonstrates a geomean speedup of $2.09\times$, $1.42\times$ and $2.83\times$ over three baselines we study (§4), respectively.
Combining Theory and Practice. One key contribution of this work is bridging theoretical analysis and empirical designs to achieve both load balance and low communication. Prior theory-driven approaches (e.g., [im2023massively, goodrich2011sorting]) achieve optimal bounds but adopt impractical techniques which lead to poor real-world performance, while empirical designs (e.g., [vora2014aspire, li2020pegasus, gog2016firmament, murray2013naiad]) lack theoretical guidance (no theoretical bounds in computation/communication costs), limiting their generalizability and robustness against skewed (or even adversarial) workloads. Our work fully redesigns theoretically optimal algorithms using practical implementation techniques such as distributed push-pull, achieving strong performance in both theory and practice. See §2.3 and §3.6 for further details of the comparison against prior work.
Case Study: Distributed Graph Processing. Among the applications listed in §2.1, graph processing is likely one of the most important. To further demonstrate the power of TD-Orch, we develop TDO-GP, a distributed graph processing system atop TD-Orch. TDO-GP leverages TD-Orch as the primary mechanism for orchestrating task-data flow and employs implementation techniques to fully exploit the task-data layouts provided by TD-Orch. Like TD-Orch, TDO-GP is theoretically grounded, offering work-efficient bounds for the five graph algorithms implemented in this study. In addition to directly leveraging TD-Orch, we further optimize global communication, local computation, and the coordination between system components, tailoring these improvements for graph workloads to fully exploit the execution layout provided by TD-Orch. As a result, TDO-GP achieves geomean speedups of $4.1\times$ over the best prior open-source distributed graph processing systems designed for general graph problems [zhu2016gemini, mofrad2020graphite, ahmad2018la3, yan2015effective] and up to $15\times$ in specific datasets.
Overall, TDO-GP outperforms two major families of prior graph systems: compared with graph-algorithm-based systems (e.g., [zhu2016gemini, chen2019powerlyra, yan2015effective]), it provides significantly better provable and empirical load balance through the use of transit machines; and compared with linear-algebra-based approaches (e.g., [mofrad2020graphite, ahmad2018la3, anderson2016graphpad]), it achieves stronger graph-theoretic guarantees and higher performance by better exploiting graph sparsity. Moreover, the interface enables user implementation of a complex graph algorithm in just 70 lines of C++ code, versus prior 400 lines in [zhu2016gemini] (see §5).
Main Contributions. In summary, in this paper:
  1. (1)
    We are the first to cleanly define task-data orchestration, a fundamental execution pattern in distributed applications, and to advocate a concrete interface for it (Fig. 1).
  2. (2)
    We design TD-Orch, a theoretically grounded and practically fast orchestration system that achieves scalable load balance with low communication overhead.
  3. (3)
    We develop TDO-GP, a distributed graph processing system built on TD-Orch and enhanced with tailored implementation techniques, which significantly outperforms prior state-of-the-art systems [zhu2016gemini, mofrad2020graphite, ahmad2018la3, yan2015effective].
2 Preliminaries
2.1 TD-Orch: Supported Applications
Many distributed applications can be abstracted using the task-data orchestration interface. We show here several examples.
Graph Processing. Many vertex- and edge-centric graph algorithms share a common execution pattern: a vertex array stores intermediate values; in each stage, a subset of edges are activated, read the values of their endpoints, perform a computation using these vertex values and the edge weight, and write the result back to the vertex array. Examples include breadth-first search (BFS), shortest paths, betweenness centrality, connectivity, PageRank, and spanning forests.
All such graph applications can be expressed within the task-data orchestration framework. Example pseudocode [^1] This example demonstrates how graph algorithms can be written using the orchestration interface. It is intentionally simple but not work-efficient. Work-efficient implementations rely on auxiliary structures such as DistVertexSubset, introduced later in §5. for BFS is shown in Algorithm˜1, where Tasks correspond to edges and Data correspond to vertex values—the distance from the source node in BFS. At each stage, lambda $f$ is applied to each edge. It checks whether the source vertex of the edge was activated in the previous round, having a distance of $\textsc{round}-1$. If so, it writes distance round to the destination vertex, provided it has not been written already.
Input: $G=(V,E)$, where $|V|=n$, $|E|=m$; StartVertex
Output: Dist[$1:n$]: Distrib. vertex value array
[…]
2 VertexUpdated = 1
3 round = 1
4 Struct Task
5 LocalContext: Edge (u, v)
6  InputPointer: pointer(dist[u])
7  OutputPointer: pointer(dist[v])
8  Function f(value_u: int) :
9     return round If (value_u == round - 1) Else -1
10Function WriteBack(dist_ptr: pointer, value_v: int) :
11  If *dist_ptr == -1 and value_v != -1 then
12     *dist_ptr = value_v
13     VertexUpdated = 1
14Function BFS($G$, StartVertex: int) :
15  Dist[StartVertex] = 0
16  Initialize Task array tasks[1:m] using G.E[1:m]
17  While VertexUpdated > 0 do
18     VertexUpdated = 0
19     Orchestration(tasks, WriteBack)
20     round += 11
21 return Dist[$1:n$]
Algorithm 1 Breadth-First Search (BFS)
Key-Value Store. Many key-value store abstractions naturally fit within the task-data orchestration model. A straightforward example is a (distributed) hash table, where reading and updating a batch of items can be expressed as a one-stage orchestration by defining $f$ as the per-item operation. Further details are provided in §6.
Similarly, many ordered indexes (e.g., B-trees, radix trees) can be represented within this abstraction by decomposing each search/update query into multiple orchestration stages. In each stage, a single level of the index/tree is traversed: the corresponding internal node is retrieved, and $f$ performs the search within the node to identify the appropriate child. This process continues until the leaf is reached.
Transactional/Analytical Databases. Transactional (OLTP) and analytical (OLAP) databases share structural similarities with key-value stores, often relying on tree-based index structures. Their index traversals can be orchestrated in the same way as key-value lookups to support search and range scan queries. In addition, OLTP/OLAP systems maintain version chains, which resemble concurrent bucket-based hash tables and are likewise supported by the orchestration framework.
Vector Search. High-dimensional approximate nearest neighbor (ANN) search is a key component in vector search. Many state-of-the-art implementations use bucket-based designs [tian2023db], graph-based approaches [jayaram2019diskann, malkov2018efficient], or hybrid schemes [munoz2019hierarchical, manohar2024parlayann] that maintain a nearest-neighbor graph at higher levels to guide searches toward candidate buckets for detailed vector comparisons. The bucket traversal naturally fits the orchestration framework, similar to a key-value store, while the graph traversal resembles a vertex-centric graph algorithm, which can also be expressed using orchestration.
Fault-Tolerant Storage. Our orchestration framework can also support RAID [chen1994raid] or erasure coding [balaji2018erasure] in fault-tolerant storage. By setting InputPointers to fetch the original data, WB to write back to the target location at OutputPointers, and $f$ to perform the RAID or erasure-coding computation, a redundancy-based erasure-resilient storage abstraction can be efficiently implemented across the machines.
2.2 System Modeling
The task-data orchestration interface maps the execution flow onto a distributed system whose details—such as scheduling and physical data partitioning—are not exposed to the user.
Model. To provide better quantified analysis on our and prior designs, we use the Bulk-Synchronous Parallel (BSP) model [valiant1990bridging, valiant1990general, valiant1988optimally]. The modeling assumes the system consists of $P$ machines, each with a processing unit and a local memory without shared global memory. All machines are connected by a point-to-point communication network. The system operates in rounds (or supersteps), where periodic barrier synchronizations are required, similar to MPI.
BSP accounts for computation time (maximum work) and communication time (maximum data exchanged) across all machines. Since the metrics use “maximum”, maintaining load balance is crucial. Readers can refer to Appendix A for more details of the model.
Data Storage. Data are partitioned into data chunks of granularity $B$ words, analogous to cache-line or page storage to exploit locality. Each chunk is placed on a random machine, providing adversary-resistant load balance in both storage and access patterns [sanders1996competitive]. Such randomized placement is commonly used to ensure robustness in high performance [wang2023provably, kang2022pimtreevldb, shan2023explore, luo2023smart].
Executed Tasks. A batch of $n$ tasks, each with a context size of at most $\sigma$ words, will be executed during an application. At initialization, the tasks are evenly distributed across the machines. During each orchestration stage, each task requires access to at most $D$ data chunks (possibly zero) stored outside its local context, and a task completes only after all required data chunks have been fetched, processed and written back.
Design Goals. Both tasks and data chunks are allowed to be moved between machines, and even routed through intermediate machines. Our goal is to orchestrate tasks and data—i.e., to design an online scheduling scheme that (i) minimizes the amounts of communication and computation, and (ii) ensures load balance across machines to maximize throughput.
Load Balance. A key challenge in achieving load balance across the machines is the contention for the data chunks, since up to $n$ tasks may require the same data chunk. We seek orchestration that is highly effective in practice and also with formal guarantees regardless of the degree of data chunk contention. For the latter, we define the following goal:
Definition 1 (Load-Balanced Stage).
A stage of tasks with total computation work $W$ and total communication $I$ is (asymptotically) load-balanced iff each of the $P$ machines incurs $O(W/P)$ computation work and $O(I/P)$ communication cost.
2.3 Scheduling Strategies
As described in §2.2, task-data orchestration basically seeks an execution scheme in which $n$ task contexts, each of size $O(\sigma)$, access $O(nD)$ data chunks, each of size $B$. The ideal communication time of $O(nD\min\{B,\sigma\}/P)$ is unachievable in practice using only online scheduling algorithms. Instead, prior work typically adopts one of the following strategies.
Direct Pull. A naïve scheduling strategy first eliminates duplicate requests for data chunks within each machine, then fetches all required chunks to the corresponding tasks. This approach, common in remote direct memory accesses (RDMA), works well when contention (reference count) on each chunk is low. However, contention can be high—for example, in graph processing, contention can be caused by extremely high-degree vertices—leading to high communication on machines storing “hot” chunks and severe load imbalance. In the worst case, all $n$ tasks may request the same chunk, incurring $O(DPB/\min\{D,P\})$ communication time and competitive ratio of $O\left(\frac{P^{2}B}{n\min\{B,\sigma\}\min\{D,P\}}\right)$, which is prohibitive as $P$ grows.
Direct Push. Another common strategy in distributed settings, such as in remote procedure calls, is to offload tasks to the machines that store the required data chunks. However, machines holding “hot” data chunks incur high communication cost (for receiving tasks) and high computation work (for executing received tasks), resulting in severe load imbalance. In the worst case, all $n$ tasks are pushed to the same chunk, resulting in $O(nD\sigma/\min\{D,P\})$ communication time and a competitive ratio of $O\left(\frac{P\sigma}{\min\{B,\sigma\}\min\{D,P\}}\right)$, which is also prohibitive.
Theory-Guided Designs. Prior work on the Massively Parallel Computing (MPC) model [im2023massively] has proposed theoretically-efficient solutions to the orchestration problem (e.g., [im2023massively, goodrich2011sorting]). The procedure begins by sorting tasks according to the addresses of their required data chunks. Then, MPC broadcasting [ghaffari2019massively] is used to distribute each data chunk to the corresponding tasks. After task execution, a reverse broadcasting step writes the results back to the data chunks, followed by a reverse sorting step that restores tasks to their original order. It has been shown that the computation work is load balanced, and the communication time is $\Theta(\frac{n}{p}\log_{\frac{n}{P}}{P})$ [fish2015computational], which is optimal [roughgarden2018suffles]. Although these MPC schemes are asymptotically optimal, we show in §4 that their constant factors are non-negligible in practice, resulting in poor performance on real-world systems (e.g., [axtmann2015practical]) compared to empirical designs.
Empirical Designs. Numerous empirical designs (e.g., [vora2014aspire, li2020pegasus, gog2016firmament, murray2013naiad]), while not directly targeting task-data orchestration, employ practically efficient techniques in their scheduling, such as communication aggregation. However, few provide rigorous analysis of competitive ratios, and none offer tight bounds, limiting their ability to generalize to broader distributed systems and robustness to adversary workloads.
2.4 Prior Distributed Graph Processing
Classic distributed frameworks such as MapReduce [dean2008mapreduce] and Pregel [malewicz2010pregel, bu2014pregelix, han2015giraph] inspired many early distributed graph systems, including BLAS [bulucc2011combinatorial], GraphLab [low2012distributed], GraphMat [sundaram2015graphmat], Graphulo [gadepally2015graphulo], GraphX [gonzalez2014graphx], Mizan [khayyat2013mizan], Pegasus [kang2009pegasus], PowerGraph [gonzalez2012powergraph], PowerSwitch [xie2015sync], and Presto [venkataraman2013presto].
We classify more recent graph systems developed in the past decade into two major families: (i) Graph-algorithm-based: GeminiGraph [zhu2016gemini], Gluon [dathathri2018gluon], PowerLyra [chen2019powerlyra], FaasGraph [liu2024faasgraph], Phoenix [dathathri2019phoenix], ShenTu [lin2018shentu], and TopoX [li2019topox]; and (ii) Linear-algebra-based: Graphite [mofrad2020graphite], GraphPad [anderson2016graphpad], GraphScope [fan2021graphscope], GraphTap [mofrad2019efficient], and LA3 [ahmad2018la3].
Compared with prior graph-algorithm-based systems, our design, TDO-GP, achieves stronger provable and empirical load balance by using transit machines in its underlying TD-Orch to migrate contention. In contrast, prior mirror/ghost-node designs (e.g., [zhu2016gemini, gonzalez2012powergraph]) alleviate only limited contention and still behave largely as direct task-data exchanges.
Compared with prior linear-algebra-based designs, TDO-GP achieves stronger theoretical bounds and higher practical performance, by better exploiting graph sparsity than sparse-matrix approaches, by using work-efficient graph algorithms.
3 Main Design: TD-Orch
TD-Orch is a practically efficient orchestration framework with provable guarantees. Aligned with the original three-phase orchestration, TD-Orch refines Phase I (Read Data) by splitting it into two subphases—contention detection and task-data co-location. This yields a four-phase execution pipeline:
  1. (1)
    Contention Detection: A preprocessing phase before the actual task-data co-location are carried out. The information of each task is sent to the machine storing its target data to gather the contention information for each data chunk. To preserve load balance, messages are passed along a communication forest network (§3.1), in the form of meta-tasks (§3.2).
  2. (2)
    Task-Data Co-location: Using the information gathered in Phase 1, tasks and data are redistributed across machines to achieve an optimized schedule that minimizes communication and computation time. (See §3.3.)
  3. (3)
    Task Execution: Now that all tasks and the copies of their target data are co-located on the same machine, the tasks are executed locally.
  4. (4)
    Data Write-backs: The updated value for any data modified by an executing task is written back to the original data location. (See §3.4.)
Key Design: Push-Pull. To address the challenges of direct push or pull, we must handle potential contention. Our method, TD-Orch, exploits the two directions of data flow: push, where a task context is sent to the machine(s) storing its required data chunks, and pull, where data chunks are fetched to the machine where the task resides. Such bidirectional strategies (referred to as distributed push-pull) are rarely emphasized in prior work. Only a few prior works demonstrate that this strategy can significantly improve load balance, but they are limited to PIM settings [kang2022pimtreevldb, kang2023pimtrie, zhao2025optimal, zhao2026pimzd, kang2025pim] and rely on an extremely powerful centralized coordinator (the host CPU). Our goal is to develop a more widely applicable push-pull technique that does not require a centralized coordinator.
Ideally, when contention information is available on any machine on the fly, distributed push-pull aggregates redundant information on one side (task/data) and dynamically chooses the flow direction toward the contended end. This dynamic decision should balance load with low communication.
3.1 Communication Forest for Handling Skew
Although the system assumes a fully connected network, tasks and data cannot be directly co-located, as this may cause severe load imbalance when hot data chunks exist. A circular challenge exists: we need a preprocessing Phase 1 to gather contention information for later load balancing, but Phase 1 itself must stay load-balanced before contention is even known.
In Phase 1 in TD-Orch, rather than exchanging all task-data in a single communication round, a tree-structured routing network in Fig. 2 is used to gradually push the information of tasks (which reside in the leaf machines and need to access a data chunk in the root machine of the tree) toward their target data across multiple rounds within Phase 1, moving up one tree level per round. This prevents tasks accessing the same hot data chunk from overloading a single machine.
Formally, a communication tree rooted at machine $i$ is a network with a balanced tree structure with $P$ leaves and fanout $F$, as shown in Fig. 2. Each of the $P$ leaf nodes represents a different physical machine, while internal nodes correspond to virtual transit machines, each of which is mapped to one of the $P$ machines via a hash function globally known to all machines. Notably, a physical machine can simultaneously serve as both a leaf and an internal node.
The communication tree is used to collect information about all tasks that need to access data chunks stored on Machine $i$. It enables efficient collection of task-related information (such as counting task occurrences) with load-balanced communication. The format of information passed along this network will be introduced in §3.2.
In the example in Fig. 2, Phase 1 needs three rounds (the height of trees). Suppose Machines 2 and 3 each have a task accessing data on Machine 1. In Round 1, each task sends a packed message (data address and task info) to its parents, VM(1,3)=PM3 and VM(1,4)=PM6. In Round 2, these messages are forwarded to VM(1,1)=PM5. In the final round, an aggregated message is delivered to Machine 1. This message aggregation prevents the overloading of a single machine.
A communication forest consists of $P$ such communication trees, each rooted at a distinct machine. All tasks send messages in parallel along the trees associated with the machines that store their target data chunks. Although the internal transit machines are statically chosen within the communication forest, each data chunk is stored at a random address (§2.2). This randomization in the corresponding tree selection is equivalent to dynamically selecting transit nodes in both theoretical analysis and practical implementation.
We use $F=\Theta\left(\log P/\log\log P\right)$ in TD-Orch based on the theoretical analysis in §3.5, which also yields strong practical performance as shown in §4and §6.
Figure 2: A communication forest with fanout $F=2$ for $P=8$ machines. Green nodes are source machines (task storage), red nodes are destination machines (data storage), and blue nodes are intermediate transit virtual machines (VMs) with IDs $(\text{root},\text{BFS\_id})$, mapped to physical machines (PMs) via example hash function $h(x,y)=(x+3y)\bmod 8+1$. Dashed edges mean that PM does not need to send messages to itself.
3.2 Meta-Task Structure
This section describes the messages passed through the communication forest during Phase 1. These messages are used to support co-location in Phase 2 and should include (i) the reference count for each data chunk, and (ii) the locations of the tasks requesting them. They should also be aggregatable to prevent unbounded message size growth at any machine.
To meet these goals, we introduce a meta-task structure—parameterized by hyperparameter $C$ and attribute LL—to serve as the message payload within the tree-structured network. At a high level, if the reference count of a data chunk is $\leq C$, the meta-task contains the full contexts of all requesting tasks. If the reference count exceeds $C$, the meta-task includes only aggregated metadata, including the count and the locations of tasks, thereby preventing unbounded message growth.
The attribute $L$, which serves as a level indicator, enables hierarchical aggregation of meta-tasks. Raw task contexts are treated as $L_{0}$ meta-tasks. During message passing in the communication forest, if an intermediate machine receives more than $C$ $L_{i}$ meta-tasks, it locally stores them and creates a new $L_{i+1}$ meta-task. This new $L_{i+1}$ meta-task contains two pieces of information: (i) the aggregated reference count, and (ii) the addresses of the machines where the $C$ underlying $L_{i}$ meta-tasks are stored. Only this aggregated $L_{i+1}$ meta-task will be forwarded to its parent machine.
A meta-task set, as shown in Figure˜3, is a set structure containing multiple meta-tasks. The number of meta-tasks in a meta-task set with the same $L$ value should not exceed $C$.
A merge operation of two meta-task sets is proposed to support message aggregation. A new meta-task set is generated, containing all meta-tasks from both of the original sets. If the number of meta-tasks at a certain level $L_{i}$ in the new set exceeds $C$, all $L_{i}$-meta-tasks will be popped out of the new set and stored locally on the current machine, and then a new $L_{l+1}$-meta-task will be pushed into the set. The merge will be executed from the lowest level up to the highest level, until the number of meta-tasks at each level is at most $C$.
Example Fig. 4 with $C=3$ shows the merge process. First, $L_{0}$-meta-tasks are merged; since their count exceeds $C$, they are locally stored and an $L_{1,\text{new}}$ is created. Next, the (now >C>C) $L_{1}$-meta-tasks, including $L_{1,\text{new}}$, are similarly stored and produce an $L_{2,\text{new}}$. There are $\leq C$ $L_{2}$-meta-tasks, so all are emitted to the output set. Continue this until the highest level.
This merging process guarantees that each meta-task set has bounded size $C\log_{C}n$, while containing sufficient location information to enable efficient task-data exchange in Phase 2.
We use $C=\Theta(B/\sigma)$ in TD-Orch following the theory-guided design in §3.5, which also leads to strong practical performance as shown in §4 and §6.
Figure 3: A meta-task set, zooming in $L_{0}$ and $L_{i}$ ($i\geq 1$) meta-tasks. Deeper colors indicate higher levels. $L_{0}$ contains original task contexts, while $L_{i}$($i\geq{1}$) contains a pointer to a $L_{i-1}$-meta-task array stored on a machine.
Figure 4: Merging two meta-task sets with $C=3$. Meta-tasks on lowest level $L_{0}$ are merged first and generate $L_{1,new}$. Then $L_{1}$-meta-tasks (including both the received $L_{1}$ and the newly generated $L_{1,new}$) are merged and generate $L_{2,new}$. Continue the process until the top level of $L_{4}$.
3.3 Distributed Push-Pull for Co-location
After Phase 1, each data chunk receives a meta-task set that contains all information needed for task-data co-location. In Phase 2, we introduce distributed push-pull, a strategy that exploits the bidirectional flow of tasks and data to resolve potential contentions on data chunks.
Push. If the reference count of a data chunk is $\leq{C}$, then all task contexts requesting this chunk have already been delivered as $L_{0}$ meta-tasks to the machine storing the chunk. Thus, for these uncontended tasks, the push process—sending tasks to data—is already completed during Phase 1.
Pull. For contended tasks, where the reference count of a data chunk exceeds $C$, tasks are stored on intermediate transit machines in the communication forest. A pull process—sending data (copies) to the tasks—is required in Phase 2. The location information in meta-task sets is used to guide the pull.
As an example, suppose a data chunk receives a meta-task set with the highest level $L_{i}$. It first broadcasts a copy of the chunk to all $\Theta(C)$ machines storing the $L_{i-1}$-meta-tasks, then chases pointers and broadcasts the data to the $\Theta(C^{2})$ machines storing the $L_{i-2}$-meta-tasks, continuing until all $L_{0}$ meta-tasks (the original task contexts) have received a copy.
This broadcasting follows a path of an inherent tree-structured network formed by a meta-task set, which we define as the meta-task tree. This tree is more compact than the communication tree, forming a full $C$-ary tree structure with all leaf machines containing $L_{0}$-meta-tasks and no empty leaves.
Key Takeaway. Bi-directional distributed push-pull is a key design that enables load-balanced execution under contention while maintaining efficient communication overheads. It combines push and pull: uncontended tasks are directly pushed to their target data, whereas contended tasks pull their data gradually, level by level, along the inherent meta-task tree.
3.4 Write Backs
If tasks require write-backs that are merge-able (defined below), they can be efficiently aggregated, allowing the write-backs from multiple tasks to the same data chunk to be combined. In Phase 4, these write-backs are propagated along the reverse path of the meta-task tree defined in §3.3, applying this aggregation process during message passing.
Definition 2 (Merge-able Operations).
An operation $\oplus$ is merge-able if there exist operations $\odot$ and $\otimes$ such that for any data $x$ and $y_{i}(1\leq i\leq n)$, $x\oplus y_{1}\oplus...\oplus y_{n}=x\odot(y_{1}\otimes...\otimes y_{n})$.
Merge-able operations $\oplus$ are commonly seen, including and not limited to (i) idempotent operations ($\odot$ is $\oplus$ and $\otimes$ is keeping one left), (ii) set-associative operations ($\odot$ and $\otimes$ are both $\oplus$), (iii) concurrent writes where only one random write operation succeeds in a concurrent batch ($\odot$ is write and $\otimes$ is randomly select), and (iv) write operations where concurrent writes are resolved using a deterministic decision process, e.g., in a concurrent batch only the one with the smallest timestamp / transaction ID succeeds ($\odot$ is write and $\otimes$ is finding the value to write using the decision process).
3.5 Formal Guarantees
We use the BSP model [valiant1990bridging, valiant1990general, valiant1988optimally] to provide theoretical guarantees for our design. Readers less interested in the theory can directly refer to the conclusions in Key Takeaway.
Our main Theorem˜1 proves TD-Orch to be asymptotically optimal compared with the lower bounds proved in [roughgarden2018suffles]. Please refer to Appendix B for its proof.
Theorem 1 (Main Theorem).
Assume $n=\Omega(\frac{PB\log^{3}{P}}{\log\log{P}})$ and $n=O(P^{\epsilon})$ for some $\epsilon>1$. If each machine holds $\Theta(n/P)$ tasks when a stage begins, and write backs are merge-able, then (i) TD-Orch takes $O(\frac{nD}{P}(\min\{B,\sigma\}+\log_{\frac{n}{P}}{P}))$ communication time whp  [^2] We use $O(f(n))$ with high probability (whp) (in $n$) to mean $O(cf(n))$ with probability at least $1-n^{-c}$ for $c\geq{1}$. in this stage; and (ii) each machine will execute $\Theta(\frac{n}{P})$ tasks whp and hold them at the end of this stage.
Key Takeaway. (i) Optimal Communication: TD-Orch incurs optimally minimal communication time as $P$ scales up. (ii) Optimal Computation: TD-Orch balances computation works, maximizing system throughput. (iii) Inductive Execution: Task storage remains load-balanced after each stage, enabling efficient execution of subsequent stages.
3.6 Comparison Against Prior Methods
Against Direct Push/Pull. TD-Orch achieves the optimal theoretical bounds as shown in §3.5, outperforming load-balance-oblivious methods.
Against Prior Theory-Guided Methods. Though asymptotically the same (both optimal), TD-Orch requires only $2$ sweeps over the communication forest, whereas MPC sorting [im2023massively, goodrich2011sorting] incurs $\geq{3}$ passes. This reduction reflects the potential of TD-Orch for strong practical efficiency.
Against Prior Empirical Designs. TD-Orch (i) provides provable guarantees, and (ii) explicitly targets the orchestration abstraction, rather than a fused design of scheduling inside specific end-to-end applications.
4 Case Study I: Key-Value Store
A natural fit for our orchestration framework is a key-value store, which we implement atop TD-Orch using a concurrent distributed hash table as the underlying storage structure.
Methods. We compare four orchestration methods: TD-Orch, direct-push, direct-pull, and sorting-based. All of them use MPI for inter-machine communication. Direct-push, direct-pull, and TD-Orch are implemented directly, while the sorting-based method is built on KaDiS [axtmann2015practical, axtmann2018lightweight, axtmann2017place, harsh2019histogram], a state-of-the-art massively parallel sorting library. All four methods follow the orchestration interface in Figure˜1.
Machine Details. Our experiments use a cluster of 16 machines, each equipped with AMD Opteron 6272 processors (64 cores, 2.1 GHz), 64 MB L3 cache, and 128 GB of memory. The machines are interconnected via a 10 Gbps network.
Workloads. We evaluate the orchestration frameworks on YCSB workloads [cooper2010benchmarking] A (50% reads, 50% writes), B (95% reads, 5% writes), C (read-only), and LOAD (write-only), with data access patterns following a Zipf distribution [zipf2016human], a common setting in database benchmarks (e.g., [kang2022pimtreevldb, luo2023smart]). Each task fetches an item from the key-value store, performs a multiply-and-add operation, and then optionally writes the updated value back.
Results. Figure˜5 shows the results of weak-scaling experiments with $P=2,4,8,16$ machines. In this setup, each machine is initialized with 2 million tasks, with target data addresses drawn from Zipf exponents $\gamma=1.5,\ 2.0,\ 2.5$.
TD-Orch is (i) resilient to adversarial skew and delivers near-linear performance under weak scaling, and (ii) significantly faster than other methods. On multi-machines, TD-Orch achieves geometrically-averaged speedups of $2.09\times$, $1.42\times$, and $2.83\times$ over the other three baselines, respectively.
Figure 5: Runtime results (in seconds) for YCSB workloads A (50% reads, 50% writes), C (read-only), and LOAD (write-only). TD-Orch scales well and outperforms prior methods. Results for YCSB-B (95% reads, 5% writes) exhibit similar trends and are omitted due to space constraints.
5 Case Study II: Graph Processing Systems
As another case study, we implement a distributed graph processing system called TDO-GP on top of our TD-Orch framework to demonstrate the effectiveness of TD-Orch. Following the common practice in prior vertex-centric works [zhu2016gemini, wu2021seastar, wang2023scaleg, zhu2022vctune], TDO-GP maintains two primary storage structures: one for vertices, which hold their intermediate values and are pinned to specific machines based on a predefined schema during graph ingestion; and another for edges and their weights, which can be dynamically orchestrated by the system.
Graph Preliminaries. We denote a graph by $G=(V,E)$ where $V$ is the vertex set and $E$ is the edge set. $n=|V|$ and $m=|E|$ are respectively the number of vertices and edges.
Interface. Our graph processing engine is primarily executed through a combined interface of TD-Orch (Figure˜1) and the EdgeMap abstraction [ShunB2013, shun2015ligraplus, dhulipala2017julienne, dhulipala18scalable].
Ligra [ShunB2013, shun2015ligraplus, dhulipala2017julienne] and GBBS [dhulipala18scalable] introduce EdgeMap, which works on a single machine and takes as input a graph $G=(V,E)$, a VertexSubset $U$ (a subset of vertices), and two boolean functions, $F$ and $C$. It applies $F$ to each edge $(u,v)\in E$ iff $u\in U$ and $C(v)=true$, returning a new VertexSubset $U^{\prime}$ consisting of all vertices $v$ for which $F(u,v)$ is applied and returns true. EdgeMap supports two execution modes—sparse and dense—ensuring that it runs in $O(\sum_{u\in U}\text{deg}(u))$ work (total computation) and $O(\log n)$ span (critical path length) on arbitrary inputs.
We refer to our new interface as DistEdgeMap, which is defined in Figure˜6. We treat $G$ as a directed graph. For undirected inputs, we represent each undirected edge $\{u,v\}$ as two directed edges $(u,v)$ and $(v,u)$. $U$ is a DistVertexSubset, a distributed version of VertexSubset which represents a subset of vertices across distributed machines. An execution function $f$ is applied to all edges $(u,v)\in E$ where $u\in U$, and returns a value. All returned values are aggregated by DistEdgeMap using a user-defined merge-able operator merge_value, and the result is write_back to $v$. DistEdgeMap outputs a DistVertexSubset formed up by the vertices whose write_back returns true.
As shown in Figure˜6, the DistEdgeMap interface closely resembles the single-machine EdgeMap, hiding the distributed system architecture and scheduling details. This is an intentional design to simplify programming for users.
function Dist_Edge_Map(
G=$(V,E)$: Directed Graph,
U: DistVertexSubset, // Current iteration
f: function, // Execution
write_back: function,
merge_value: Merge-able operation
) -> DistVertexSubset; // For next iteration
Figure 6: DistEdgeMap function
Example Graph Algorithms. We implement five representative graph algorithms using the DistEdgeMap in our system: breadth-first search (BFS), single-source shortest path (SSSP), betweenness centrality (BC), connected components (CC), and PageRank iteration (PR).
In all these algorithms, users only need to define the functions $f$ and write_back using only a few lines of code; then launch DistEdgeMap without needing to know the details of the underlying execution flow. Please refer to Appendix C for an implementation of a complex algorithm such as BC in fewer than 70 lines of code using our interface. This is much more concise than prior frameworks (e.g., Gemini [zhu2016gemini] requires over 400 lines).
5.1 Main Design of TDO-GP
Ingestion-Time Orchestration of Task-Data Flow. A distributed graph algorithm that can be expressed using DistEdgeMap can also be viewed as a task-data orchestration, whereas edges correspond to tasks and vertex intermediate values correspond to data (see §2.1). A key observation is that, regardless of the starting vertex, in each DistEdgeMap, the contention for data (vertex value) access induced by tasks (edges) is proportional to the degree of each vertex. Since vertex degrees are intrinsic to the structure of the graph, such contention is inevitable in all DistEdgeMap operations, regardless of the choice of starting vertex in an algorithm.
Given this observation, TD-Orch can be applied once as a preprocessing phase when the graph is ingested, which is sufficient to resolve all future skew in the execution of arbitrary DistEdgeMap operations. In future stages, DistEdgeMap leverages these pre-orchestrated flows rather than dynamically re-orchestrating at every stage. This one-time load balancing is particularly beneficial for highly skewed graphs, such as the commonly encountered power-law graphs [gonzalez2012powergraph].
The preprocessing process consists of two stages. In the first stage, we begin by randomly placing each edge on an arbitrary machine and then execute one round of TD-Orch, which attempts to colocate all edges with their source vertices. Edges whose sources are low-degree vertices are now placed on the same machine with the vertices. For edges with high-degree sources, a tree structure in meta-tasks (denoted as source trees) is created to represent their source-side communication. We then fix the storage location of these edges and propagate future source-vertex values through the source trees. In the second stage, another round of TD-Orch is performed, where each edge, starting from its location at the end of the first stage, tries to reach its destination vertex. This again results in a tree structure in meta-tasks (called destination trees), this time representing the paths along which the values to the destination vertices will be write_backwrite\_back.
Adopting this preprocessing provides two key benefits: (i) it organizes edge storage into chunks (possibly on transit machines, rather than on source/destination machines, if highly contended), ensuring balanced computation load; and (ii) it establishes a tree-shaped communication framework for both get_value and write_back, enabling aggregation communication. This design differs from prior ghost/mirror node designs (e.g., [zhu2016gemini, gonzalez2012powergraph]), which neither use transit machines to mitigate contention nor employ tree-structured networks to improve communication costs. As a result, our design achieves better load balance and improved performance (§6).
Sparse-Dense Execution. We design a dual-mode execution flow for distributed graph processing which can switch between sparse and dense modes to efficiently exploit the data layouts provided by TD-Orch. The mode is determined by the size of the current DistVertexSubset. Specifically, the sparse mode is optimized for small DistVertexSubset, while the dense mode is optimized for large ones.
In the sparse mode, a vertex-centric execution flow is used: each active vertex $u\in U$ propagates its value down the source tree, activating all corresponding edges. The result values are then sent and merged along the destination tree.
In contrast, the dense mode uses an edge-centric execution flow. All active vertices $u\in U$ broadcast their values to all machines. Each edge $(u,v)$ is then executed if $u\in U$.
The choice between sparse and dense mode is dynamically determined by the size of the active frontier in the current round. Specifically, if $\Sigma_{u\in{U}}deg(u)=o(P\cdot|U|)$ where $d(\cdot)$ denotes the out-degree of inter-machine edges, the sparse mode is used; otherwise, the dense mode is used. This dynamic switching ensures that the execution remains both work- and communication-efficient in theory (with work and communication of $O(\Sigma_{u\in{U}}deg(u))$ as in Ligra and load-balanced due to TD-Orch), while also achieving high practical performance.
5.2 Implementation Techniques
In TDO-GP, we employ multiple implementation techniques, which can be grouped into three categories. Due to space limit, we summarize here the key functionalities of each group. Details are provided in Appendix D.
(T1) Optimized Global Communication. We use deduplication and destination-aware broadcasting to further reduce inter-machine communication.
(T2) Work-Efficient Local Computation. Our design leverages ParlayLib [blelloch2020parlaylib], a state-of-the-art parallel library, to implement work-efficient computations. We enhance VertexSubset with a concurrent hash table + bitmap, and further develop a lightweight local EdgeMap. Moreover, DistEdgeMap accepts an optional filter_dst function, excluding vertices that are guaranteed not to be destinations in the current round.
(T3) Aligned Coordination between Components. We design tighter coordination between system components—between threads and scheduler, among threads to prevent cache thrashing, and between local and global execution.
5.3 Formal Guarantees
Since Theorem˜1 applies to DistEdgeMap, all of our algorithms are work-efficient and load-balanced across machines. The theoretical bounds summarized in Table˜1 are achieved when $n=\Omega(P\log^{2}{P})$. While prior works [zhu2016gemini, mofrad2020graphite, ahmad2018la3, yan2015effective, lu2014large, yan2014pregel] mostly focus on empirical system improvements, we place strong emphasis on graph-theoretic guarantees as guidance for implementation. As shown in §6.2, these guarantees directly translate into practical performance improvements.
Table 1: Theoretical bounds of graph algorithms.
Alg. BFS SSSP BC $C$ PR
Comp. $O\left(\frac{m}{P}\right)$ $O\left(\frac{m}{P}\right)$ $O\left(\frac{m}{P}\right)$ $O\left(\frac{m}{P}\cdot{\text{diam}}\right)$ $O\left(\frac{m}{P}\right)$
Comm. All equal to $O(\text{Computation}\cdot{\log_{\frac{n}{P}}{P}})$
6 Evaluation of Graph Processing Systems
6.1 Experimental Setup
Machine Details. We evaluate all methods on the same 16-machine cluster described in §4, each equipped with AMD Opteron 6272 processors (64 cores, 2.1 GHz), 64 MB L3 cache, 128 GB memory, and a 10 Gbps interconnect network.
Competitors. Due to their state-of-the-art performance, we select GeminiGraph [zhu2016gemini], Graphite [mofrad2020graphite], and LA3 [ahmad2018la3] as baselines. GeminiGraph is the fastest graph-algorithm-based system, and the two are the fastest linear-algebra-based systems, representing the strongest alternatives in both design families. Other earlier classic designs, such as PowerGraph [gonzalez2012powergraph], GraphX [gonzalez2014graphx], PowerLyra [chen2019powerlyra], GraphPad [anderson2016graphpad], and Giraph [apache2017giraph], are not directly compared due to their reported lower performance than that of the three baseline methods [ahmad2018la3, mofrad2020graphite, zhu2016gemini] included in our evaluation.
Additionally, we include Pregel+ [yan2015effective, lu2014large, yan2014pregel] in one end-to-end experiment in §6.2, and the state-of-the-art single-machine design GBBS [dhulipala18scalable] in the ablation study in §6.5.
Datasets. We evaluate on Reddit [hamilton2017inductive], uk-2005 [BoVWFI, BRSLLP], Twitter-2010 [kwak2010twitter, boldi2011layered], Friendster [yang2012defining], Hyperlink-2012 [meusel2015graph], and Road-USA [road], covering a wide range of dataset sizes and graph characteristics. See the first column of Table˜2.
6.2 End-to-End Results
Table˜2 presents a performance comparison between our proposed framework, TDO-GP, and three state-of-the-art systems: Gemini [zhu2016gemini], Graphite [mofrad2020graphite], and LA3 [ahmad2018la3], evaluated on various graph datasets. We select the number of machines roughly proportional to the number of vertices in each dataset, except for particularly demanding datasets (e.g., Road-usa) that require excessive runtime. Overall, TDO-GP consistently achieves superior performance, outperforming the baselines in 28 out of 30 tested scenarios, most by substantial margins.
For social media graphs such as Twitter-2010, Friendster, and Reddit, TDO-GP achieves speedups more than $2\times$ in $19$ out of the $30$ tests we carry out. These gains stem from the ability of TDO-GP to (i) more effectively exploit sparsity (sparse-dense execution in §5.1 and techniques in §5.2), and (ii) minimize inter-machine communication while achieving better load balance (ingestion-time TD-Orch in §5.1).
For large-diameter graph datasets, such as Hyperlink-2012 (web graph), uk-2005 (web graph), and Road-USA (road network), the performance gains of TDO-GP are even more pronounced. The most striking example is Road-USA, where TDO-GP achieves more than $15\times$ speedup over Gemini and over $100\times$ compared to Graphite and LA3. These improvements arise from the complexity of $O(n\cdot\text{diam}+m)$ work bounds in the baseline algorithms, compared with the work bounds of $O(n+m)$ in TDO-GP. In high-diameter graphs, where few vertices are active per round but the number of rounds is large, baseline inefficiencies are amplified, whereas TDO-GP maintains stable and efficient performance.
TDO-GP underperforms the baselines in only $2$ cases out of the $30$ experiments: Twitter PR and Road-USA PR. Detailed explanations for these exceptions will be provided in §6.5.
We also compared TDO-GP against Pregel+ [yan2015effective, lu2014large, yan2014pregel]. However, their Hadoop-based implementations are at least $5\times$ slower than all the other baselines considered here and are therefore omitted from subsequent analyses.
Table 2: End-to-end runtime comparison (in seconds) across various graph datasets and algorithms.
Dataset Alg. TDO-GP Gemini Graphite LA3
Reddit BFS 0.015 0.151 0.130 0.157
4 machines SSSP 0.071 0.172 0.141 0.505
n=2.33n=2.33M BC 0.076 0.275 0.507 0.437
m=114m=114M $C$ 0.016 0.211 0.266 0.663
diam [^2] Computing the exact diameter of large graphs is often prohibitively expensive. In our results, an asterisk (*) denotes an estimated diameter obtained using the approximation method in Ligra [ShunB2013], whereas diameters without an asterisk correspond to values reported by the dataset providers. $6^{*}$ PR 0.143 0.676 0.297 1.92
uk-2005 BFS 0.759 8.98 8.01 32.56
8 machines SSSP 0.820 10.68 13.79 45.41
n=39.5n=39.5M BC 3.00 20.23 34.79 75.90
m=482m=482M $C$ 3.30 11.22 20.85 63.11
diam $276^{*}$ PR 2.37 2.77 7.69 9.06
Twitter2010 BFS 0.938 1.35 3.13 7.67
8 machines SSSP 2.05 3.63 3.36 7.94
n=41.7n=41.7M BC 3.04 4.34 6.16 15.54
m=1.47m=1.47B $C$ 1.28 7.65 7.40 21.73
diam $23^{*}$ PR 10.46 12.72 9.85 13.67
friendster BFS 1.98 3.12 3.70 12.38
8 machines SSSP 3.09 4.42 4.25 27.23
n=65.6n=65.6M BC 5.29 8.24 7.81 69.15
m=1.80m=1.80B $C$ 2.52 17.92 9.92 56.25
diam =32=32 PR 18.32 33.07 21.71 28.03
Hyperlink12 BFS 1.60 6.22 5.59 20.50
16 machines SSSP 2.40 8.02 28.53 33.10
n=102n=102M BC 3.03 16.09 46.15 51.50
m=0.93m=0.93B $C$ 5.26 11.71 29.42 48.31
diam $95^{*}$ PR 7.12 10.10 14.43 38.78
Road-USA BFS 10.68 357.73 1224.83 1431.11
16 machines SSSP 15.76 362.11 1651.52 1788.53
n=23.9n=23.9M BC 26.24 763.09 >>35min >>35min
m=28.9m=28.9M $C$ 12.93 383.51 1814.93 >>35min
diam $6139^{*}$ PR 1.00 0.612 3.01 3.59
6.3 Scalability
Strong Scaling. We run the SSSP and BC algorithms on the Twitter dataset using 1, 2, 4, 8, and 16 machines. The results in Figure˜7 demonstrate that TDO-GP exhibits near linear scalability across all machine configurations, while other methods either show poor scalability when increasing the number of machines, or exhibit poor performance.
Weak Scaling. We also evaluate weak scaling on the graph systems by varying the number of machines while keeping the graph size per machine fixed. To construct graphs on such a scale, we employ two graph generation methods: Erdős–Rényi (ER) [erd6s1960evolution], which produces unskewed graphs, and Barabási–Albert (BA) [barabasi1999emergence], which produces skewed graphs. The average number of edges per machine is fixed at 40M for a fair comparison, where in the 16-machine case the graph is of a similar size as uk-2005. For the BA model, we set the power-law exponent to $\gamma=2.2$, consistent with the measured skew observed in natural graphs reported by PowerGraph [gonzalez2012powergraph]. We run both PR and BC on $P=1,2,4,8,16$ machines. In Figure˜9, the x-axis represents the number of machines and the y-axis is the execution time. The results show that TDO-GP maintains nearly constant runtime across machine counts, demonstrating excellent weak-scaling performance, whereas prior works exhibit significant scalability degradation.
(a) SSSP on Twitter.
(b) BC on Twitter.
Figure 7: Strong scaling results on Twitter dataset with 1–16 machines. TDO-GP is significantly faster and more scalable.
(a) PR on ER Graph
(b) PR on BA Graph
(c) BC on ER Graph
(d) BC on BA Graph
Figure 8: Weak scaling results. $y$-axis is runtime (in seconds) and $x$-axis is the number of machines used. TDO-GP exhibits excellent weak-scaling performance and is significantly faster.
6.4 Ablation Study
Execution Time Breakdown. As shown in Figure˜9, on 16 machines, the time spent on pure local computation is effectively reduced across all algorithms. At the same time, the communication time (mainly for propagating vertex values and the cost of synchronization rounds) and the overhead time (including data preparation and data receiving for multi-machine transmission) are also non-negligible, and together contribute significantly to the overall execution time.
Sensitivity to Each Optimization. We first demonstrate the effectiveness of TD-Orch within TDO-GP by comparing its final design against its prototype of Ligra plus direct-pull, which does not use TD-Orch. As shown in Table˜3, our BC results on the Twitter graph achieve up to $220\times$ speedup.
Additionally, we evaluate the sensitivity of TDO-GP to the implementation techniques introduced in §5.2. Using a controlled-variable approach, we measure the slowdown incurred when removing each optimization family from the fully optimized TDO-GP. Table˜4 reports the results on Twitter with $P=4,8,16$, showing that disabling any single optimization family causes substantial performance degradation, with slowdowns of up to $5.72\times$, illustrating the importance of each technique in achieving high practical efficiency.
Figure 9: Runtime on Twitter with 16 machines. Breakdown into Communication, Computation, and Overhead.
Table 3: Runtime (in seconds) of BC on Twitter shows that TD-Orch is a critical building block in TDO-GP.
Number of Machines 1 4 8 16
Ligra-Dist (no TD-Orch) 5.36 563.11 670.71 321.60
TDO-GP 4.54 3.70 3.05 2.35
Table 4: Slowdown when removing individual techniques (see §5.2) from the fully optimized TDO-GP.
Number of Machines 4 8 16
Final Optimized TDO-GP (Runtime in seconds).
BFS: TDO-GP 1.04 0.94 0.87
SSSP: TDO-GP 2.65 2.06 1.54
BC: TDO-GP 3.70 3.05 2.35
PR: TDO-GP 15.80 11.91 9.03
Removing (T1): Optimized Global Communication.
[…]
Removing (T2): Work-Efficient Local Computation.
[…]
Removing (T3): Aligned Coordination.
[…]
6.5 Ablation on PageRank and NUMA Effects
As shown in Table˜2, the only 2 out of the 30 cases where TDO-GP underperforms prior systems are both PageRank (PR), and only against one prior baseline. TDO-GP is built upon ParlayLib [blelloch2020parlaylib], a fast parallel library that achieves good scalability with respect to the total number of cores, but is NUMA-oblivious. While this design works well on machines with high-speed NUMA interconnects, our 16-machine cluster has budget constraints: each machine contains four NUMA nodes connected in a square topology, where certain NUMA-to-NUMA accesses are significantly slower. This architectural limitation hinders the performance of TDO-GP on PR.
We conduct two additional ablation studies to evaluate the effectiveness of TDO-GP when NUMA-related limitations are mitigated. First, we perform a multi-machine experiment on the same cluster used in §6.2, but restrict execution to a single NUMA node per machine. The results, shown in Table˜5, demonstrate that the performance gap between TDO-GP and baselines is substantially reduced. Moreover, TDO-GP has a significantly smaller memory footprint than Graphite. The remaining minor performance difference arises because ParlayLib [blelloch2020parlaylib], the underlying parallel library of TDO-GP, is optimized for many-core settings, making the single-NUMA configuration less favorable.
We further conduct a single-machine experiment on a more powerful server with superior NUMA interconnects. This machine is equipped with four Intel(R) Xeon(R) E7-8867 v4 @2.40GHz processors, providing a total of 144 cores and 180MB of L3 cache. Unlike our 16-machine cluster, it features an all-to-all interconnection across its four NUMA nodes, eliminating the bottlenecks of square-topology interconnects. The results on the Twitter dataset are presented in Table˜6, where TDO-GP consistently outperforms all other baselines by a significant margin. Notably, TDO-GP even surpasses GBBS [dhulipala18scalable], the state-of-the-art single-machine implementation, owing to our optimized lightweight local EdgeMap.
Given the results of these ablation studies, we are confident that TDO-GP can be efficiently deployed to address all the graph problems mentioned in this paper on clusters with fast NUMA interconnections, which are typical in high-performance computing environments and datacenters.
Table 5: Runtime (in seconds) of PR on Twitter. Each machine uses only 11 NUMA node. OOM is out-of-memory.
Machine Num. 1 4 8 16
Gemini 46.67 30.72 25.61 18.59
Graphite OOM OOM 22.18 14.69
TDO-GP 40.74 32.30 21.67 15.07
Table 6: Twitter runtime (in seconds) on Xeon E7-8867 v4.
Method BFS BC PR
Gemini 0.29 1.17 4.72
Graphite 2.17 3.91 10.05
GBBS 0.29 1.05 6.81
TDO-GP 0.12 0.95 4.43
7 Conclusion
In this paper, we introduce an abstraction of the task-data orchestration interface, unifying a wide range of distributed applications. We design TD-Orch to realize this orchestration with scalable load balance and low communication. Building on TD-Orch, we develop TDO-GP, a distributed graph processing system enhanced with implementation techniques for graph workloads. Both TD-Orch and TDO-GP are theoretically grounded and achieve high practical performance.
Acknowledgements
This research was supported by the Parallel Data Lab (PDL) Consortium (Amazon, Bloomberg, Datadog, Google, Honda, Intel, Jane Street, LayerZero, Meta, Microsoft, Oracle, Pure Storage, Salesforce, Samsung, and Western Digital).
Appendix A Models for Distributed Computing
In this paper, we model and analyze algorithms using the classic Bulk-Synchronous Parallel (BSP) model [valiant1990bridging, valiant1990general, valiant1988optimally]. The BSP consists of $P$ machines, each with a sequential processing element (PE) and a local memory. There is no shared global memory. All machines are connected by a point-to-point communication network. All remote accesses to data on other machines are carried out via messages sent over this network. The system operates in bulk-synchronous rounds (supersteps), where periodic barrier synchronizations separate periods of asynchronous computation and communication.
Cost Metrics. The BSP model defines the notion of an $h$-relation of a superstep, which combines the maximum computation work, maximum communication amount and synchronization time over all machines. Due to using maximum when defining $h$-relations, load balance is critical for achieving good performance. In all our analysis, for clarity, we decouple the computation time and communication time, and do not consider the BSP parameters $g$ and $L$ in [valiant1990bridging, gerbessiotis1994direct]. Computation time $t$ and communication time $h$ can be combined into a formal BSP result $\Theta(gh+t)$ as long as the synchronization overhead can be hidden by system communication, i.e., $(gh+t)\in\Omega(L)$.
Other Distributed Models. The Massively Parallel Model of Computation (MPC) [im2023massively, karloff2010model, goodrich2011sorting, beame2013communication, andoni2014parallel] family of models regards synchronization as the main bottleneck in distributed systems and takes communication rounds as the only optimization target. We choose the the BSP model because it differs from MPC in two ways: (i) The the BSP model accounts for the total communication among machines rather than ignoring the amount of communication in each round. (ii) The the BSP model also considers computation costs neglected by the MPC model, which turn out to be critical for real-world system designs on emerging distributed hardware [kang2022pimtreevldb, bindschaedler2020hailstorm] other than MapReduce systems  [dean2008mapreduce].
LOCAL/CONGEST models [linial1987distributive, linial1992locality] are a popular model family for distributed systems. Like the models we consider, each machine has only a local view of the computation. These models consider network topology, which we ignore in the BSP model. However, they have similar weaknesses to the MPC model with regards to their metrics: (i) each machine can send/receive up to its maximum allowed communication at each round, whereas we seek to minimize communication, and (ii) computation costs are not considered.
The Processing-in-Memory (PIM) model [kang2021processing] is a recently proposed model for emerging compute-in-memory hardware [mutlu2023primer]. The model has a powerful Host CPU and $P$ PIM nodes, each comprised of a PE and a memory. It operates in bulk-synchronous rounds, but unlike the BSP model, all communication must be sent by or to the Host CPU. Furthermore, the model decouples the computation time of the CPU from other processors, resulting in a lack of analysis of load-balance between them.
Appendix B Proof for Theorem˜1
We will use the balls-into-bins game of Lemma˜1 to prove load balance.
Lemma 1 (Weighted Balls into Bins  [sanders1996competitive, nievergelt1973binary]).
Placing weighted balls with total weight $W=\Sigma w_{i}$ and weight limit $W/(P\log P)$ into $P$ bins uniformly randomly yields $\Theta(W/P)$ weight in each bin whp.
Proof for Theorem˜1.
When $n=\Omega(P^{\epsilon})$ for some $\epsilon>1$, things become trivial by setting $F=\Theta(n/P)$. Here we only analyze the hard case when $n=O(P\log^{\epsilon}{P})$ for some $\epsilon>1$. We set $C=\Theta(B/\sigma)$ and $F=\Theta(\log{P}/\log{\log{P}})$. (i) Phase 1 takes at most $O(\log_{F}{P})$ rounds and in each round at most $O(n)$ words of messages are sent. Since meta-task sizes are bounded by $O(C\log_{C}{n})\leq{O}\left(\frac{n}{P\log{P}}\right)$, the weight limit in Lemma˜1 is met and the communication is thus load-balanced. When $n=\Omega\left(\frac{PB\log^{3}{P}}{\log\log{P}}\right)$, the communication time is $O(\frac{n}{P}\log_{F}{P})=O(\frac{n}{P}\log_{\frac{n}{P}}{P})$ in phase 1. Phase 2 takes less communication time than phase 1 because it only requires partial communication over phase 1’s trace, so (i) is proved. (ii) Since meta-tasks have size $O(C\log_{C}{n})$, Lemma˜1 can also be applied. ∎
Appendix C Example Algorithms for Graph Problems
Algorithm˜1 presents pseudocode for breadth-first search (BFS), offering a more efficient implementation via DistEdgeMap than directly using the orchestration interface.
In addition, Algorithm˜3 is presented here to describe the pseudocodes for the betweenness centrality (BC) computation from a single root vertex, which is commonly used in performance tests in prior works [ShunB2013, dhulipala18scalable, zhu2016gemini].
We also provide the actual C++ implementation of the BC algorithm in Figure˜10 using the DistEdgeMap interface. Despite the complexity of the original algorithm, the code remains highly concise, requiring fewer than 70 lines.
Appendix D Details of Implementation Techniques
We present here the implementation techniques used in our graph processing system, which efficiently exploit the data layout and execution flow provided by TD-Orch, leading to the strong performance demonstrated in §6.
D.1 Optimized Global Communication
As we will show in §6.4, the critical bottleneck in a distributed graph system is the inter-machine communication time. Optimizing global communication is therefore essential. First, we apply deduplication during message passing along the meta-task trees, aggressively merging all repeated items. Second, in dense mode, we reduce broadcast traffic by sending the values of each vertex only to the machines that store its corresponding edges. The mapping of vertices to machines is established during the preprocessing phase.
D.2 Work-Efficient Local Computation
For all local computations in our implementation, we employ work-efficient designs that are both theoretically optimal and practically fast. Our design leverages ParlayLib [blelloch2020parlaylib], a state-of-the-art C++ parallel library providing a wide range of work-efficient parallel algorithms as primitives.
The DistVertexSubset structure consists of a collection of VertexSubset instances, one per machine, where each VertexSubset can independently switch between two representation formats introduced in Ligra [ShunB2013]. We introduce two implementation improvements. (i) In the sparse representation, we replace their array of vertices with a phase-concurrent hash table with adaptive sizing, enabling more efficient lookups within VertexSubset. (ii) In the dense representation, we replace the original parallel C++ Boolean-map with a concurrent bitmap, improving cache efficiency.
For local edges whose source and destination vertices reside on the same machine, we bypass the global orchestration framework and instead execute a local EdgeMap. We implement a lightweight EdgeMap entirely using ParlayLib primitives, which, as we show in §6.5, is more efficient than the ones in Ligra/GBBS for EdgeMap workloads.
Additionally, we allow users to provide an optional input function filter_dst to DistEdgeMap to accelerate execution by filtering out vertices that are guaranteed not to be destination vertices in the current round. Disabling filter_dst does not affect correctness but may reduce efficiency.
D.3 Aligned Inter-Component Coordination
A distributed graph processing system involves multiple software components running across diverse hardware resources. Ensuring a coordinated execution schema among these components is crucial. Our implementation optimizes this coordination to achieve aligned and efficient execution.
For better alignment with the parallel thread scheduler in ParlayLib, we avoid using nested ParallelFor, which is commonly seen in graph algorithms but inefficient under such schedulers. Instead, we adopt Flattern, assisted by delayed structures [blelloch2020parlaylib], to transform two-dimensional nested parallelism into one-dimensional parallelism.
To better align local execution across threads and minimize cache thrashing, we design concurrent counters and other shared concurrent structures with sizes padded to a multiple of the cache-line size. This padding prevents false sharing, ensuring that accesses from different cores do not contend on the same cache line.
We also align local computation with the load-balanced data layout across machines. We initialize the graph with a vertex-value layout such that the total number of outgoing edges (out-degrees) assigned to each machine is approximately equal. This ensures that most local computations are naturally load-balanced across machines, since each handles a comparable number of edges.
Input: $G=(V,E)$, where $|V|=n$; StartVertex
Output: Dist[$1:n$]: Distrib. vertex value array
[…]
2 round=1=1
3 Function f ( (u,v): Edge ) :
4  return round IF (Dist[$u$] == round - 1) ELSE -1
5Function write_back ( v: Vertex, aggr_value ) :
6  If Dist[$v$] == -1 and aggr_value != -1 then
7     Dist[$v$] = aggr_value
8     return True
9 return False
10Function BFS ( $G$, StartVertex ) :
11  DistVertexSubset frontier = {StartVertex}
12  Dist[StartVertex] = 0
13  While frontier is not empty do
14     frontier = DistEdgeMap ($G$, frontier, f, write_back, MAX)
15     round += 11
16 return Dist[$1:n$]
Algorithm 2 Breadth-First Search (BFS)
Input: $G=(V,E)$; StartVertex
Output: BCValues: distributed vertex value array
1 Initialize distributed vertex value array: NumPaths[$1:n$], BCValues[$1:n$], rounds[$1:n$]
2 round=1=1
3 Function f_forward ( (u,v): Edge ) :
4 return NumPaths[$u$]
5Function write_back_forward ( v: Vertex, aggr_value ) :
6 If rounds[$v$] == 0 or rounds[$v$] == round then
7    NumPaths[$v$].fetch_add(aggr_value)
8    If rounds[$v$] == 0 then
9       rounds[$v$] = round
10       return True
11 return False
12Function f_backward ( (u,v): Edge ) :
13 return BCValues[$u$]
14Function write_back_backward ( v: Vertex, aggr_value ) :
15 If rounds[$v$] == round then
16    BCValues[$v$].fetch_add(aggr_value)
17 return False
18Function BC ( $G$, StartVertex ) :
19 DistVertexSubset frontier = {StartVertex}
20 frontiers = {frontier}
21 NumPaths[StartVertex] = 1
22 rounds[StartVertex] = 1
 // Forward Pass
23 While frontier is not empty do
24    round += 11
25    frontier = DistEdgeMap ($G$, frontier, f_forward, wb_forward, add)
26    frontiers.append(frontier)
 // Backward Pass
27 parallel for $i\in[1:n]$ do
28    BCValues[$i$] = NumPaths[$i$] = 1/1/NumPaths[$i$]
29 While round >0>0 do
30    frontier = frontiers[round]
31    DistEdgeMap ($G$, frontier, f_backward, wb_backward, add)
32    round -= 11
33
34 parallel for $i\in[1:n]$ do
35    BCValues[$i$] = BCValues[$i$] / NumPaths[$i$] - 1
36 return BCValues[$1:n$]
Algorithm 3 Betweenness Centrality (BC)
Figure 10: The actual C++ implementation of betweenness centrality (BC) using our DistEdgeMap interface is highly concise—fewer than 70 lines of code.